### v1

# Beyond Statistical Estimation: Differentially Private Individual Computation in the Shuffle Model

Abstract
The shuffle model of differential privacy (DP) has recently emerged as a powerful one for decentralized computation without fully trustable parties. Since it anonymizes and permutes messages from clients through a shuffler, the privacy can be amplified and utility can be improved. However, the shuffling procedure in turn restricts its applications only to statistical tasks that are permutation-invariant.
This work explores the feasibility of shuffle privacy amplification for prevalent non-statistical computations: spatial crowdsourcing, combinatorial optimization, location-based social systems, and federated learning with incentives, which suffer either computationally intractability or intolerable utility loss in existing approaches (e.g., secure MPC and local DP). We proposes a new paradigm of shuffle model that can provide critical security functionalities like message authorization and result access control pertained to permutation-equivariant tasks, meanwhile maintaining the most of privacy amplification effects. It incurs almost the same computation/communication costs on the client side as the non-private setting, and permits the server to run arbitrary algorithms on (noisy) client information in plaintext. Our novel technique is introducing statistically random identity into DP and force identical random distribution on all clients, so as to support secure functionalities even after message shuffling and to maintain privacy amplification simultaneously. Given that existing DP randomizers fails in the new shuffle model, we also propose a new mechanism and prove its optimality therein. Experimental results on spatial crowdsourcing, location-based social system, and federated learning with incentives, show that our paradigm and mechanism is fast as non-private settings, while reducing up to $90\%$ error and increasing utility performance indicates by $100\%$-$300\%$ relatively, and can be practical under reasonable privacy budget.
Index Terms:
differential privacy, shuffle model, digital identities, spatial crowdsourcing, federated learning
I Introduction
Personal information fuels a wide array of data-driven applications, e.g. statistical analytics, machine learning, recommendation systems, spatial crowdsourcing, mobile crowdsensing, combinatorial optimization, and location-based social systems. These applications deliver substantial value but depend on data collected from users, which is a prime target for attacks and carries a high risk of leakage or abuse. Data privacy concerns are escalating, especially after several high-profile data breach incidents. Despite the introduction of stricter privacy laws such as the EU’s General Data Protection Regulation, the California Consumer Privacy Act, and China’s Personal Information Protection Law, many users still distrust service providers and are hesitant to consent to the use of their data. To bridge this trust gap and encourage user participation, significant efforts have been made recently to develop privacy-enhancing technologies that enable private data processing without relying on a trusted third party.
Two prominent technologies addressing this problem are secure multiparty computation (MPC) [88] and differential privacy (DP) [28]. MPC uses interactive cryptographic protocols to allow mutually untrusted parties to jointly evaluate a function on their private data, with each party receiving an output but nothing else. However, the cryptographic nature of MPC imposes heavy overheads, making it challenging to scale for practical use. DP, on the other hand, adds random noise to data to ensure privacy, making it computationally more feasible since the computation is carried out on sanitized data without heavy cryptographic machinery. Traditional central DP still requires a trusted curator [28], but the local model (LDP [53]) allows each user to sanitize their data locally before sending it out. Due to its practicality and low trust assumption, LDP has been adopted by companies like Apple [74], Microsoft [24], and Google [30] in their real-world systems. However, LDP’s drawback is that each user must independently inject sufficient random noise to guarantee privacy, leading to significant utility loss.
Recently, within the realm of DP research, the shuffle model [17, 29] has emerged. This model introduces a shuffler that randomly permutes messages from users and then sends these anonymized messages to the computing server or analyzer. The server can then compute on these messages to derive the result. Trust in the shuffler is minimized because each user encrypts their locally sanitized data using the server’s public key. This way, the shuffler remains oblivious to the messages received from the users. A key advantage of the shuffle model over LDP is utility. It has been proven that anonymizing/shuffling messages amplifies the privacy guarantee provided by the local randomizer used by the users. For instance, shuffled messages from $n$ users each adopting local $\epsilon$-DP actually preserve differential privacy at the level $\epsilon_{c}=\tilde{O}(\sqrt{e^{\epsilon}/n})$ [31, 32]. Consequently, to achieve a predefined global privacy goal, each user consumes less privacy budget when sanitizing their data locally, meaning less noise needs to be added. This significantly improves the accuracy of the final result. Owing to these utility advantages, extensive studies have been conducted within the shuffle model, e.g., [10, 37, 38, 38, 39].
That said, the shuffle model has a noticeable limitation. The privacy amplification effect relies heavily on anonymizing/shuffling messages, which significantly restricts the types of computation that can be performed. So far, the sole form of computation achievable within the shuffle model is statistical estimation, i.e., the server takes the shuffled messages, aggregates them, and computes a single output from them, e.g., a count, sum, or histogram. However, many real-world applications are non-statistical in nature. When multiple users pool their data together for joint computation, they expect an individualized output that may differ for each user. We coined the term “individual computation” for such tasks. Examples of individual computation tasks include:
  • I.
    Combinatorial optimization: Spatial crowdsourcing [80], advertisement allocation [61], and general combinatorial optimization [55], where two or more parties are often matched together based on their private information. Each party should get their own list of “best matches” whatever that means.
  • II.
    Information retrieval: Mobile search [52], location-based systems [4, 22], where the query results (e.g., nearby restaurants or neighboring users) depend on the private information of the inquirer.
  • III.
    Incentive mechanisms: In federated learning [89] or data crowdsensing [72], incentives play a vital role to encourage well-behaved participation. The amount of rewarding incentive must be computed for individual users based on their contribution (e.g., via Shapley values [69]).
At first glance, private individual computation seems unattainable within the shuffle model because the need for individualized output clashes with the anonymization required for privacy amplification. However, this is not necessarily true. Our observation is that many individual computing functionalities are equivariant to shuffling. That is, the permutation applied to the inputs does not affect the computation. This means shuffling does not prevent the server from producing personalized answers for each client—the server does not know and does not need to know which answer is for whom. Nevertheless, there is a problem: how to pass the output back to the right user without breaking anonymity. A straightforward approach is to let the shuffler maintain a long-term duplex connection channel between each client and the server (e.g., as in an Onion routing network [41]), but this method is costly (due to storing communication states) and may be subject to de-anonymization attacks on anonymous channels [63, 65]. Additionally, the duplex connection is not supported in current shuffler implementations, such as the one in the seminal work [17]. This is the first technical challenge we need to address.
The second challenge we face is how to design optimal randomizers for individual computation in the shuffle model. Often the randomizer in a shuffle model protocol needs to be tailored to the tasks. For statistical tasks within the shuffle model, several studies have developed near-optimal randomizers, as seen in histogram estimation [36, 31] and one-dimensional summation estimation [9]. However, the new setting of individual computation is different: the focus is on the accuracy of the output for each user (instead of the statistical accuracy of the population). This gap renders existing randomization strategies (e.g., randomizers utilizing dimension sampling, budget splitting, or data sketching, as reviewed in [86]), along with prevalent randomizers (such as adding Laplace [28] or Staircase [35] noises), less effective for the new setting. Therefore we need to take a fresh look at the fundamental privacy-utility trade-offs and re-design the underlying randomizers.
In this work, we introduce a new paradigm of the shuffle model that provides security functionalities (e.g., message authorization, integrity verification, and access control) for permutation-equivariant computations, and term it the private individual computation model. The properties of the model are as follows.
  • I.
    Security: The server and other clients can verify the owner of a message using a fingerprint submitted alongside the message. The server can also securely deliver personalized backward information to each owner (i.e., access control). These functionalities enable the secure completion of equivariant tasks.
  • II.
    Privacy: There is no degradation of local privacy while providing these additional security functionalities. Furthermore, when task results are sparse (e.g., each user is matched to one taxi driver in spatial crowdsourcing), the degradation of shuffle privacy amplification is negligible. Additionally, our local and shuffle privacy guarantees are information-theoretic (as long as the shuffling is statistically random). This distinguishes our new model from computational DP [62], which also combines cryptographic functions and privacy but guarantees differential privacy only against polynomial-power adversaries.
  • III.
    Utility: Since clients still benefit from privacy amplification via shuffling, the utility is significantly improved. With the increasing of number of clients, the utility increases and approaches to non-private settings. In addition, the server in our model can run arbitrary algorithms on the (noisy) messages in plaintext. Our model is compatible with existing achievements and future advancements in server-side algorithms (e.g., noisy-aware matching algorithms for spatial crowdsourcing).
  • IV.
    Scalability: As the server runs algorithms on plain-text messages and there is no heavy computation burden on the client side, our model is highly efficient and scales to systems with massive clients. This eliminates the need for secure multi-party computation which necessitates impractical costs. In our model, the computation/communication costs are essentially the same to non-private settings (see Appendix -I).
Main techniques. Our approach to realizing additional security functionalities is based on two critical observations: (1) If the message payload satisfies $\epsilon$-LDP, appending arbitrary random variables (independent to the private value) to the payload incurs no additional privacy loss. Specifically, when random variables are sampled from an identical distribution for all clients, there is no excess loss in privacy amplification via shuffling. (2) Post-processing the message payload and appended random variables does not incur excess privacy loss. The first observation motivates us to append a random fingerprint (e.g., a public key) to the payload, allowing the client to use the corresponding secret key to negotiate shared secret and establish secure communication with other parties in later phases. A side consequence is that the server can use the public key pertained to each message to securely deliver information to the message owner. The second observation motivates us to further append a signature to protect the integrity of the fingerprint and payload.
Challenge (2): Optimal randomizers for non-statistical computations in the shuffle model. In the local model of DP, existing research [83, 66, 77, 81, 82] primarily focuses on designing randomizers in a high privacy regime (e.g., local budget $\epsilon=O(1)$). Conversely, in the shuffle model, the local budget can extend to $\tilde{O}(\log(n))$. For statistical tasks within this model, several studies have developed near-optimal randomizers, as seen in histogram estimation [36, 31] and one-dimensional summation estimation [9]. However, the new setting of non-statistical computation tasks in the shuffle model—often involving multi-dimensional user data and emphasizing an individual report’s accuracy (instead of statistical accuracy of the population)—has not been explored. This gap renders existing randomization strategies (e.g., randomizers utilizing dimension sampling, budget splitting, or data sketching, as reviewed in [86]), along with prevalent randomizers (such as adding Laplace [28] or Staircase [35] noises), less effective for non-statistical computation in the shuffle model. The fundamental privacy-utility trade-offs in non-statistical computation for this new model also present significant theoretical interest.
Our work establishes the error lower bounds for single reports in non-statistical computation within the shuffle model, demonstrating the sub-optimality of existing randomizers. To address this, we introduce a novel series of local randomizers—the Minkowski response—aligned with these error lower bounds. The core principle involves reporting values within a calibrated Minkowski distance from the true value with high probability.
The main contributions of this work are summarized as follows:
  • We propose equivariant shuffle model of DP, which supports various security functionalities and benefit large-scale equivariant computations from privacy amplification. The new model exhibits almost the same computation/communication costs on client/server side as non-private settings.
  • We establish tight minimax error lower bounds of $\tilde{O}((n\epsilon^{2}_{c})^{-\frac{2}{d+2}})$ for the new model of DP, where $d$ is the user data dimension, $\epsilon_{c}$ is global privacy budget, and $n$ is the number of users. Since existing data randomizers fail, we then design an attainable randomizer: Minkowski response, for the new model to match the lower bound.
  • Through experiments on three typical permutation-equivariant tasks: privacy-preserving spatial crowdsourcing, location-based nearest neighbor queries, and federated learning with incentives, we demonstrate that the new model, in conjunction with proposed randomizers, reduces error by $90\%$ meanwhile incurring negligible excess burdens in both the client and server side.
Organization. The remainder of this paper is organized as follows. Section II reviews related works. Section III provides preliminary knowledge about privacy definitions and security primitives. Section IV formalizes the problem setting. Section V presents the PIC model. Section VI provides optimal LDP randomizers. Section VII evaluates the utility and efficiency performances of our proposals. Finally, Section VIII concludes the paper.
II Related Works
This section reviews various approaches to private computation, primarily concentrating on permutation-equivariant tasks.
II-A Cryptography Methods
Numerous works aim to securely solve matching and allocation problems involving personal sensitive information from multiple parties, revealing only the output. They utilize homomorphic encryption [2, 44], secret sharing [3] or garbled circuits (e.g., in [40, 15]) to perform secure multi-party computations (MPC). Specifically, when there is a trusted message shuffler, every client can establish information-theoretically secure communication with other clients and accomplish arbitrary multi-party computation (including DP algorithms) within a few rounds [48, 13]. In general, MPC-based approaches impose high computation and interaction overheads, especially when the computation algorithm gets sophisticated. As comparison, the proposed paradigm in this work permits running arbitrary algorithms on plaintext, and requires only semi-trustness/honest-but-curious assumption on the shuffler (as the shuffler only sees ciphertexts, see Section V).
II-B Curator and Local DP Methods
Many works study matching, allocation, or general combinatorial optimization problems within the curator DP model [28] in the presence of a trusted party collecting raw data from clients (e.g., in [60, 23, 76, 75]). Since the assumption of a trustworthy party is often unrealistic in decentralized settings, many studies adopt the local model of DP (e.g., in [83, 66, 77, 81]), where each client sanitizes data locally and sends the noisy data to the server for executing corresponding matching/allocation algorithms. As each client must injects sufficient noises into data to satisfy local DP, the execution results often maintain low utility.
II-C Shuffle Model of DP
The recently proposed shuffle model [17, 29] combines the advantages of the curator model (e.g., high utility) and the local model (e.g., minimal trust). Depending on the number of messages each client can send to the intermediate shuffler, the shuffle model can be categorized as single-message [9, 29, 32] and multi-message [10, 36]. The single-message shuffle model leverages privacy amplification via shuffling to enhance data utility compared to the local model. A substantial body of work [29, 31, 32] demonstrates that $n$ shuffled messages from clients, each adopting a same $\epsilon$-LDP randomizer, can actually preserve $\tilde{O}(\sqrt{e^{\epsilon}/n})$-DP. By removing the constraint of sending one message, the multi-message shuffle model can achieve better utility than the single-message model and might be comparable to the curator model (e.g., in [10, 36]). However, each multi-message protocol is tailored to a specific statistical query (e.g., summation), rendering them unsuitable for permutation-equivariant tasks with non-linear computations. There is a line of works on the shuffle model for private information retrieval (e.g., in [48, 47, 34] with cryptography security and in [78, 1] with statistical DP), where the query is represented as multiple secret shares before sent to the shuffler, and the server holding the database entries returns linear-transformed entries for each query, using the duplex shuffled communication channel. This kind of duplex-communication shuffle model can be vulnerable to anonymity attacks [63, 65], and is pertained to the linear computation in private information retrieval. It can not be extended to other PIC tasks that involve with non-linear computations, and can not handle party-to-party communication in tasks like spatial crowdsourcing and social systems.
Overall, existing works in the shuffle model primarily focus on statistical queries that are naturally permutation-invariant. This work, for the first time, explore the shuffle model for non-statistical applications (i.e., combinatorial optimization, location-based social systems, and incentive mechanisms). Technically, our work extends the shuffle model to support identity-related security primitives, such as message authentication and access control.
II-D Combining Cryptography and DP
While cryptographic tools can protect data secrecy during multi-party computation, they do not necessarily preserve output’s privacy. DP can be employed to enhance the privacy of the outputting result of secure multiparty computation through decentralized noisy addition [43]. To account for privacy loss due to intermediate encrypted views in MPC, researchers have proposed the relaxed notion of computational DP [62] against polynomial-time adversaries. Computational DP protocols inherited the computation/communication complexity of MPC (refer to garbled-circuits-based approaches in Table VIII).
In contrast to enhancing relaxed privacy in the cryptographic cipher domain, our work initializes the study in the opposite direction, by integrating simple security functionalities into the privacy realm. Our model exposes intermediate results with information-theoretical DP guarantee. Therefore, it is significantly more efficient (i.e., without computation on ciphertext and requiring few rounds of interaction) and flexible (i.e., supporting arbitrary computations on noisy plaintext) than MPC with computational DP.
III Background and Problem Settings
TABLE I: List of notations.
Notation Description
$[i]$$\{1,2,...,i\}$
$[i:j]$$\{i,i+1,...,j\}$
$\mathcal{S}$ the shuffling procedure
$\mathcal{R}$ the randomization algorithm
$G_{i}$ the $i$-th group of users ($i\in[m]$)
$n_{i}$ the number of users in group $G_{i}$
$u_{i,j}$ the $j$-th user in group $G_{i}$ where $j\in[n_{i}]$
$d$ the dimension of user’s data
$\mathbb{X}$ the domain of input data
$\mathbb{Y}$ the domain of a sanitized message
$\epsilon$ the local privacy budget
$\epsilon_{c}$ the amplified privacy level
$sk$, $pk$ the secret key and the public key, respectively
$\lambda$ the security parameter of cryptography
III-A Privacy Definitions
Definition III.1 (Hockey-stick divergence [70])
For two probability distributions $P$ and $Q$, the Hockey-stick divergence between them with parameter $e^{\epsilon}$ is as follows:
$D_{\epsilon}(P||Q)=\int_{z\in\mathcal{Z}}\max\{0,P(z)-e^{\epsilon}Q(z)\}\mathrm{d}z.$
Differential privacy imposes divergence constraints on output probability distributions with respect to changes in the input. In the curator model of differential privacy, a trusted party collects raw data $x_{i}\in\mathbb{X}$ from all users to form a dataset $T=\{x_{1},\ldots,x_{n}\}$ and applies a randomization algorithm $\mathcal{R}$ to release query results $\mathcal{R}(T)$. For two datasets $T$ and $T^{\prime}$ of the same size and differing in only one element, they are referred to as neighboring datasets. Differential privacy ensures that the Hockey-stick divergence between $\mathcal{R}(T)$ and $\mathcal{R}(T^{\prime})$ is bounded by a sufficiently small value (i.e., $\delta=O(1/n)$). The formal definition of curator differential privacy is as follows:
Definition III.2 ($(\epsilon,\delta)$-DP [28])
A randomization mechanism $\mathcal{R}$ satisfies $(\epsilon,\delta)$-differential privacy iff $\mathcal{R}(T)$ and $\mathcal{R}(T^{\prime})$ are $(\epsilon,\delta)$-indistinguishable for any neighboring datasets $T,T^{\prime}\in\mathbb{X}^{n}$. That is, $\max(D_{\epsilon}(\mathcal{R}(T)\|\mathcal{R}(T^{\prime})),D_{\epsilon}(\mathcal{R}(T^{\prime})\|\mathcal{R}(T)))\leq\delta$.
In the local model of differential privacy, each user applies a randomization mechanism $\mathcal{R}$ to their own data $x_{i}$, with the objective of ensuring that the Hockey-stick divergence between $\mathcal{R}(x)$ and $\mathcal{R}(x^{\prime})$ is 00 for any $x,x^{\prime}\in\mathbb{X}$ (see Definition III.3).
Definition III.3 (local $\epsilon$-DP [53])
A randomization mechanism $\mathcal{R}$ satisfies local $\epsilon$-DP iff $D_{\epsilon}(\mathcal{R}(x)||\mathcal{R}(x^{\prime}))=0$ for any $x,x^{\prime}\in\mathbb{X}$.
The data processing inequality is considered a key feature of distance measures used for evaluating privacy. It asserts that the privacy guarantee cannot be weakened by further analysis of a private mechanism’s output.
Definition III.4 (Data processing inequality)
A distance measure $D:\Delta(\mathcal{S})\times\Delta(\mathcal{S})\to[0,\infty]$ on the space of probability distributions satisfies the data processing inequality if, for all distributions $P$ and $Q$ in $\Delta(\mathcal{S})$ and for all (possibly randomized) functions $g:\mathcal{S}\to\mathcal{S^{\prime}}$,
$D(g(P)\|g(Q))\leq D(P\|Q).$
III-B The Classical Shuffle Model
Following the conventions of the randomize-then-shuffle model [21, 9], we define a single-message shuffle protocol $\mathcal{P}$ as a list of algorithms $\mathcal{P}=(\mathcal{R},\mathcal{A})$, where $\mathcal{R}:\mathbb{X}\to\mathbb{Y}$ is local randomizer on client side, and $\mathcal{A}:\mathbb{Y}^{n}\to\mathbb{Z}$ is the analyzer on the server side. We refer to $\mathbb{Y}$ as the protocol’s message space and $\mathbb{Z}$ as the output space. The overall protocol implements a mechanism $\mathcal{P}:\mathbb{X}^{n}\to\mathbb{Z}$ as follows. User $i$ holds a data record $x_{i}$ and a local randomizer $\mathcal{R}$, then computes a message $y_{i}=\mathcal{R}(x_{i})$. The messages $y_{1},\ldots,y_{n}$ are shuffled and submitted to the analyzer. We denote the random shuffling step as $\mathcal{S}(y_{1},\ldots,y_{n})$, where $\mathcal{S}:\mathbb{Y}^{n}\to\mathbb{Y}^{n}$ is a shuffler that applies a uniform-random permutation to its inputs. In summary, the output of $\mathcal{P}(x_{1},\ldots,x_{n})$ is represented by $\mathcal{A}\circ\mathcal{S}\circ\mathcal{R}(X)=\mathcal{A}(\mathcal{S}(\mathcal{R}(x_{1}),\ldots,\mathcal{R}(x_{n})))$.
In particular, when all users adopt an identical $\epsilon$-LDP mechanism $\mathcal{R}$, recent works [31, 32] have derived that $n$ shuffled $\epsilon_{l}$-LDP messages satisfy $(O((1-e^{-\epsilon})\sqrt{e^{\epsilon}\log(1/\delta)/n}),\delta)$-DP. We denote the amplified privacy level as:
$\epsilon_{c}=\texttt{Amplify}(\epsilon,\delta,n),$
the tight value of which can also be numerically computed [56, 85].
III-C Security Primitives
We employ several well-established security primitives for secure point-to-point communication, including key aggreement, symmetry encryption and digital signature.
A Diffie-Hellman key agreement protocol with security parameter $\lambda$ comprises two primary steps:
  • $sk_{i},pk_{i}\leftarrow\textsf{DH-Gen}(\lambda)$: a party $i$ generates a public-private key pair $(sk_{i},pk_{i})$, where $pk_{i}$ is the public key and $sk_{i}$ is the secret/private key;
  • $K_{ij}\leftarrow\textsf{DH-Compute}(sk_{i},pk_{j})$: the two parties involved in the key exchange use their private keys $sk_{i}$ and $sk_{j}$ to compute a shared key $K_{ij}$ from the public keys $pk_{j}$ and $pk_{i}$ of each party (i.e., $K_{ij}=K_{ji}$.
The resulting shared secret key, $K_{ij}$, can be used for subsequent data symmetry encryption/decryption. Specifically, a symmetry encryption scheme with secret key $sk$ consists of two steps:
  • $c\leftarrow\textsf{Enc}(sk,x)$: generates a ciphertext $c$ by encrypting data $x$ with the encryption key $sk$;
  • $x\leftarrow\textsf{Dec}(sk,c)$: decrypts ciphertext $c$ with the key $sk$, outputting $x$.
A digital signature scheme permits the receipt to authorize ciphertexts and to ensure message integrity, through the following steps:
  • $t\leftarrow\textsf{Sign}(sk,c)$: returns a tag $t$ as the output of signing data $c$ with the private key $sk$;
  • $accept/reject\leftarrow\textsf{Verify}(pub,t,c)$: outputs $accept$ when $t=\textsf{Sign}(sk,c)$, and outputs $reject$ with a probability of $1-negl(k)$ when $t\neq\textsf{Sign}(sk,c)$.
IV Problem Settings
We aim to extend the capability of the shuffle model to support private individual computation. This would have multiple benefits: (1) we can get a formal privacy guarantee (in the DP sense) on a much wider range of data-driven computational tasks; (2) because the computation is done on already sanitized data, we can avoid heavy cryptographic protocols and scale to large data and/or user size; (3) the privacy amplification effect will allow much better utility, compared to sanitizing and computing directly using LDP. In this section, we will set the stage by first presenting a few motivating applications, and then formally defining private individual computation as an ideal functionality.
Figure 1: An illustration of conceptual overview of PIC model for spatial crowdsourcing.
IV-A Motivating Applications
We describe three prevalent exemplar computation tasks: spatial crowdsourcing, location-based services, and federated learning with incentives:
Spatial crowdsourcing. A spatial crowdsourcing system normally consists of three roles: users, workers, and the orchestrating server (the service platform), and proceed as following major steps:
  • I.
    Task submission: each participated user $i\in G_{a}$ submits a task (e.g., taxi calling requests, sensing/photo requests) $x_{i}=(i,l_{i},v_{i})$ containing the location $l_{i}$ and possibly other information $v_{i}$;
  • II.
    Worker reporting: every enrolled worker $j\in G_{b}$ reports $x_{j}$ containing the location $l_{j}$ and other information $v_{j}$;
  • III.
    Task assignment: the server receives $\{x_{i}\}_{i\in G_{a}}$ and $\{x_{j}\}_{k\in G_{b}}$, and outputs a matching $M:G_{a}\times G_{b}\mapsto\{0,1\}$ between $G_{a}$ and $G_{b}$ based on some criterion (e.g., minimizing total traveling costs, or maximizing matches).
  • IV.
    Task performing: users and workers retrieve matching results $M$ and collaboratively complete the task.
Location-based services. In a location-based service, a set of users and a server interact as the following:
  • I.
    Querying: Each participating user $i\in G_{a}$ submits a request $x_{i}=(i,l_{i},v_{i})$, which includes information like location $l_{i}$ and preferences $v_{i}$.
  • II.
    Generating recommendation: The server receives $\{x_{i}\}_{i\in G_{a}}$ and generates a list of recommendations for each user based on specific criteria, such as proximity and user preferences.
  • III.
    Retrieving: Finally, each user retrieves the recommendation results.
Federated learning with incentives. Federated learning involves a set of clients and a server:
  • I.
    Submitting gradient: Each client $i\in G_{a}$ in each epoch computes an intermediate gradient information $x_{i}\in[-1,1]^{d}$ with its local model and data, then submits to the server;
  • II.
    Computing incentive: The server computes the averaged gradient $\overline{x}=\frac{1}{n}\sum_{i\in[n]}x_{i}$. To incentive participation, the server may reward clients with monetary tokens (e.g., via cryptocurrency) according to a profit allocation algorithm $V:[-1,1]^{d\times n}\times[-1,1]^{d}\mapsto[0,1]$ (e.g., Shapley value [69]).
  • II.
    Receiving incentive: Each client retrieves the token and claims its monetary incentive.
A fundamental distinction between the above applications and those currently studied in the shuffle model is that now each participant expects an output that differs individually, instead of a single collectively aggregated output. Informally, in such applications, we would need to safeguard data privacy so that apart from the party who generates the data, no one should be certain about that party’s data (up to the leakage allowed by DP). To be able to amplify privacy, we also need to maintain anonymity so that for any message, an adversary (e.g. the server, an observer, an unmatched user) should only know this message comes from a user belonging to a particular group, but nothing more. Note that following the convention in the shuffle model, the shuffler is trusted for providing anonymity. So the shuffler will know the random permutation used in the shuffling process and will not leak it to other parties, although we prevent the shuffler from observing plaintext messages through encryption.
IV-B The Ideal Functionality
Following the ideal-real world paradigm, we capture private individual computation formally as an ideal functionality $\mathcal{F}_{\scriptscriptstyle\textsf{PIC}}$, which is shown in Figure 2.
Functionality $\mathcal{F}_{\scriptscriptstyle\textsf{PIC}}$ Parameters: $m\in\mathbb{N}$; $m$ groups of parties $G_{1},G_{2},\cdots,G_{m}$, where $n_{i}=|G_{i}|$ denotes the number of parties in group $G_{i}$; the data randomization mechanisms $\mathcal{R}_{i}$ for group $G_{i}$; a server ${S}$. Functionality: Upon receiving $n=\sum_{i\in[m]}n_{i}$ inputs $\{x_{i,j}\}_{i\in[m],j\in[|G_{i}|]}$ from all clients, and the description of a function $f$ to be computed over parties’ inputs from the server, do the following: Compute $x^{\prime}_{i,j}\leftarrow R_{i}(x_{i,j})$ for all $i\in[m]$ and $j\in[|G_{i}|]$. Sample $m$ random permutations $\pi_{1},\pi_{2},\cdots,\pi_{m}$, where $\pi_{i}:[n_{i}]\rightarrow[n_{i}]$, perform shuffle over inputs and obtain $L=\left(\{x_{1,\pi_{1}(j)}^{\prime}\}_{j\in[n_{1}]},\cdots,x_{m,\pi_{m}(j)}^{\prime}\}_{j\in[n_{m}]}\right)$, and compute ($\left(\{y_{1,\pi_{1}(j)}\}_{j\in[n_{1}]},\cdots,\{y_{m,\pi_{m}(j)}\}_{j\in[n_{m}]}\right)\leftarrow f(L)$. Send $y_{i,\pi_{i}(j)}$ to party $u_{i,j}$ for all $i\in[m]$ and $j\in n_{i}$. Additionally, send $(L,f(L))$ to $S$.
Figure 2: The functionality $\mathcal{F}_{\scriptscriptstyle\textsf{PIC}}$
Essentially, the ideal functionality represents a fully trusted party that interacts with $m$ groups of clients and one server $S$. It sanitizes the input from each client, then shuffles them randomly, and applies a function from the server to compute the output for each client. At the end, each client receives its function output, and the server gets the list of shuffled sanitized data from the clients and a list of the function output. This captures the functional requirement of individual computation in the real world: a computing task is performed by a server using the joint inputs from a set of clients (which are sanitized and shuffled), and each client gets an individualized output. It also captures the privacy requirements: each party should receive precisely the specified output, nothing more.
Remark A careful reader may notice that the ideal function does not capture differential privacy. This is intentional for the sake of security analysis. Later in the security analysis, we will decouple different privacy requirements into two sets of proofs: the first set shows that our concrete protocol executed by real-world parties realizes the ideal functionality, meaning that no additional information about parties’ inputs is revealed, except the output given to each party; then we prove that the output conforms with differential privacy.
V A Concrete Protocol
V-A The Protocol
We now present a concrete protocol for private individual computation. Because there are already secure protocols for shuffling, to simplify the description, we present the protocol in the $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$-hybrid model, in which parties can communicate as usual, and in addition have access to an ideal functionality $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$ that does the shuffling. The ideal functionality $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$ (Figure 3) is parameterized with a list of parties that are corrupted by the adversary and collude with the server. For those parties, the adversary should know the correspondence between their messages before and after shuffling, hence $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$ leaks this part of the permutation to the adversary.
Functionality $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$ Parameters: $n\in\mathbb{N}$; $n$ parties $P_{1},P_{2},\cdots,P_{n}$; a server ${S}$; the corrupted party set $\mathcal{C}$; the leakage $\mathcal{L}(\pi)=\{(i,\pi(i))\}_{i\in[\mathcal{C}]}$ for the permutation $\pi$ being used. Functionality: Upon receiving $n$ inputs $\{x_{i}\}_{i\in[n]}$ from $P_{1},P_{2},\cdots,P_{n}$, respectively. Sample a random permutation $\pi\in\mathbf{S}_{n}$. Define $\{y_{i}\}_{i\in[n]}$ such that $y_{i}=x_{\pi(i)}$. Send $\{y_{i}\}_{i\in[n]}$ to the server $S$. Additionally if $S\in\mathcal{C}$, send $\mathcal{L}(\pi)$ to the adversary $S$.
Figure 3: The functionality $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$
The Private Individual Computation protocol is outlined below:
  1. 1.
[…]
  2. 2.
    We denote the $j$-th client in group $G_{i}$ as $u_{i,j}$. Each client generates a key pair $(pk_{i,j},sk_{i,j})\leftarrow\textsf{Gen}(\lambda)$. Each client then randomizes its private data $x_{i,j}$ with mechanism $\mathcal{R}_{i}$ and obtains $x_{i,j}^{\prime}\leftarrow\mathcal{R}_{i}(x_{i,j})$. Then the sanitized input is concatenated with its own public key, and encrypted with the server’s public key $x_{i,j}^{\prime\prime}\leftarrow\textsf{Enc}_{pk_{c}}(pk_{i,j}||x^{\prime}_{i,j})$.
  3. 3.
    Each client in group $G_{i}$ invokes $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$ with $x_{i,j}^{\prime\prime}$, and $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$ outputs the shuffled messages $\{x_{i,\pi_{i}(j)}^{\prime\prime}\}_{j\in[n_{i}]}\leftarrow\mathcal{S}(\{x_{i,j}^{\prime\prime}\}_{j\in[n_{i}]})$ to the server, where $\pi^{-1}_{i}$ is the (secret) random permutation used during $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$ for group $G_{i}$.
  4. 4.
    The server decrypts each set of shuffled messages and obtains a list $L$ for all $m$ groups as:
    $L=$ ($\left(\{pk_{1,\pi_{1}(j)}||x_{1,\pi_{1}(j)}^{\prime}\}_{j\in[n_{1}]},\cdots\right.$
    $\ \ \ \left.\{pk_{m,\pi_{m}(j)}||x_{m,\pi_{m}(j)}^{\prime}\}_{j\in[n_{m}]}\right)$
    It then computes the function $f$ over $L$ to produce output for each anonymous client:
    ($\left(\{y_{1,\pi_{1}(j)}\}_{j\in[n_{1}]},\cdots,\{y_{m,\pi_{m}(j)}\}_{j\in[n_{m}]}\right)\leftarrow f(L)$
  5. 5.
    The server publishes the computation results to a public bulletin board as a list of pairs: $\{(pk_{i,\pi_{i}(j)},\textsf{Enc}_{pk_{i,\pi_{i}(j)}}(y_{i,\pi_{i}(j)})\}_{j\in[n_{i}]}$, for each group $i\in[m]$.
  6. 6.
    Every client downloads the list, finds in the list the entry with its own public key, and decrypts the payload to get the computation result.
V-B Security Analysis
This section presents the formal security analysis of the protocol in the previous section. In the following, we will consider a semi-honest adversary who can statically corrupt parties in the protocol. That is, the adversary will faithfully follow the protocol specifications but try to learn more information than allowed through protocol interaction. Also, before running the protocol, the adversary specifies the corrupted parties. The adversary controls the corrupted parties and knows their internal states. We use $\mathcal{C}$ to denote the collection of corrupted parties and $\mathcal{C}\subset G_{1}\cup G_{2}\cdots G_{n}\cup\{S\}$.
We first show that our protocol securely realizes the ideal functionality $\mathcal{F}_{\scriptscriptstyle\textsf{PIC}}$ in the $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$-hybrid model. This means the protocol leaks no more information than what is allowed by $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$ and $\mathcal{F}_{\scriptscriptstyle\textsf{PIC}}$. More precisely, each client gets their output from $\mathcal{F}_{\scriptscriptstyle\textsf{PIC}}$, the server gets $(L,f(L))$, and in the case of colluding with some clients in $G_{i}$, the partial permutation $\mathcal{L}(\pi_{i})$. Formally we have the following theorem:
Theorem V.1 (Security)
The Private Individual Computation (PIC) protocol in §V securely computes $\mathcal{F}_{\scriptscriptstyle\textsf{PIC}}$ in the $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$-hybrid model in the presence of any PPT adversary with static corruption.
The proof can be found in Appendix -J. The proof is simulation-based. It shows that for any PPT adversary $\mathcal{A}$ in the real world, there exists a PPT simulator $\mathcal{S}$ in the ideal world that can generate a simulated view given the corrupted parties’ inputs and outputs. Security means that the simulated view is indistinguishable from the view of real-world execution.
The above theorem states that the adversary learns strictly no more than the allowed output and leakage by engaging in the protocol execution. Next, we will show how much differential privacy we can get in the presence of such an adversary with such knowledge. We consider an honest user $u_{\scriptscriptstyle i^{*},j^{*}}$, where $i^{*}\in[m],j^{*}\in[n_{i^{*}}]$. At the same time, the set of corrupted users in $G_{i^{*}}$ is denoted as $C_{i^{*}}\subset G_{i^{*}}$. Differential privacy in our case means that on two neighboring inputs $X=\big{(}X_{1},\cdots,X_{i^{*}}\cdots,X_{m}\big{)}$ and $X^{\prime}=\big{(}X_{1},\cdots,X^{\prime}_{i^{*}}\cdots,X_{m}\big{)}$, the output and leakage obtained by the adversary, denoted as $\mathcal{A}(X)$ and $\mathcal{A}(X^{\prime})$, are close in distribution. Formally, we have the following theorem:
Theorem V.2 (Differential Privacy)
The Private Individual Computation protocol satisfies $(\epsilon_{c},\delta)$-DP, i.e.
$\max(D_{\epsilon_{c}}(\mathcal{A}(X)||\mathcal{A}(X^{\prime})),D_{\epsilon_{c}}(\mathcal{A}(X^{\prime})||\mathcal{A}(X)))\leq\delta.$
In particular, when all users in $G_{i^{*}}$ use an identical $\epsilon$-LDP mechanism as the data randomizer $\mathcal{R}_{i^{*}}$, and for $n^{\prime}_{i^{*}}=|G_{i^{*}}-C_{i^{*}}|\geq\tilde{\Omega}(e^{\epsilon})$ we have:
$\small\epsilon_{c}=\log\Bigg{(}1+\frac{e^{\epsilon}-1}{e^{\epsilon}+1}\Big{(}\sqrt{\frac{32(e^{\epsilon}+1)\log 4/\delta}{n^{\prime}_{i^{*}}}}+\frac{4(e^{\epsilon}+1)}{n^{\prime}_{i^{*}}}\Big{)}\Bigg{)}.\normalsize$ (1)
The analysis can be divided into two cases: in the first case, $S\in\mathcal{C}$, i.e. S is corrupted. In this case, since the honest client locally randomizes their input, the input enjoys at least $\epsilon$-DP. Then the shuffling will amplify the privacy guarantee. Since the corrupted server receives the partial permutation as the leakage, the amplification depends on the number of uncorrupted users in the same group ($n^{\prime}_{i^{*}}$) as the honest user. The amplified $\epsilon_{c}$ can then be derived following [31, 32]. In the second case where $S\not\in\mathcal{C}$, the knowledge of the adversary is $y_{i,j}$ for each corrupted user $u_{i,j}\in\mathcal{C}$. The tricky part is that how much information $y_{i,j}$ leaks depends on the function $f$ being evaluated by the server. Hence we consider the worst case where $f$ output $y_{i,j}=L$. Continuing along the same line of thought, we conclude that the level of differential privacy assurance is no less than in the first case. The full proof can be found in Appendix -A.
V-C Discussion on Post-computation Communication
In certain scenarios, such as spatial crowdsourcing and location-based social systems, there may be additional client-to-client communications following the execution of the PIC protocol. For instance, consider a ridesharing application where passengers are in a group $G_{1}$ and taxi drivers in a group $G_{2}$. After receiving the matching result at the end of the protocol, the passengers must send their locations to the matched drivers, who need to know where to pick them up. On the other hand, the drivers also need to share their identities and locations with the matched passengers. Inevitably, a party has to sacrifice their privacy to the matched parties.
The post-computation communication may also have privacy implications for other clients not in the matched pair. Privacy amplification via shuffling against an adversary relies on the number of parties that remain anonymous. Recall that in Equation 1, the amplified $\epsilon_{c}$ relies on the number $n^{\prime}_{i^{*}}=|G_{i^{*}}-C_{i^{*}}|$ of uncorrupted users in a particular group $G_{i^{*}}$. From the perspective of the adversary, if the post-computation communication compromises the anonymity of an additional set $U$ of the clients in $G_{i^{*}}$, then $n^{\prime}_{i^{*}}$ becomes $|G_{i^{*}}-C_{i^{*}}-U|$. Accordingly, the privacy amplification effect for users in $G_{i^{*}}$ that are still anonymous is weakened.
We emphasize that preventing the loss or weakening of privacy via technical means is not feasible, because the information is necessary for the proper functioning of the application. However, managerial countermeasures, such as ensuring sufficiently large user groups and limiting post-computation exposure according to the need-to-know principle, can mitigate potential privacy risks arising from post-computation communication.
VI Optimal Randomizers
VI-A Inadequacy of Existing Randomizers
In the PIC model, each user first sanitizes their data using an $\epsilon$-LDP randomizer. The design of the randomizer significantly impacts the utility of the tasks. While the PIC model can be seen as an extension of the shuffle model, it has unique characteristics that render the existing LDP randomizers commonly used in the shuffle model inadequate.
The main discrepancy between the shuffle model and the PIC model is that the former emphasizes statistical utility, whereas the latter focuses on the utility of each report. The shuffle model aims to estimate certain statistics from the noisy data collected from users, so it cares about how close the estimation is to the true value of the desired statistic. In the literature, utility is often measured by the expected square error (i.e., the variance) of the estimation:
$\mathbb{E}_{\scriptscriptstyle T\in\mathbb{X}^{n}}[\|\tilde{f}(T)-f(T)\|_{2}^{2}]=\textsf{Var}_{\scriptscriptstyle T\in\mathbb{X}^{n}}[\tilde{f}(T)]$
where $f$ is a statistical function, and $\tilde{f}$ is its estimation output by the shuffle protocol. On the other hand, in the PIC model, the tasks are often non-statistical. For example, in location-based matching, the required computation is to take two users’ locations and compute the distance between them. Hence, the above utility measure is no longer suitable. It is more natural to measure utility by the single report error:
$\max_{x_{j}\in\mathbb{X}}\mathbb{E}[\|\mathcal{R}(x_{j})-x_{j}\|_{2}^{2}],$
where $\mathcal{R}$ is the LDP-randomizer employed by the users. Additionally, in the PIC model, user data is typically multi-dimensional (e.g., location, gradient), making sanitization significantly more difficult compared to scalar data. Another notable characteristic of the PIC model is that the local differential privacy budget $\epsilon$ is relatively large, often scaling linearly with $\log(n^{\prime}_{i^{*}})$, as implied by Theorem V.2. These factors together create issues when existing LDP randomizers are applied directly in the PIC model.
To see the problem, we first examine a class of LDP randomizers [64, 84, 58, 50]) that work by sampling a few dimensions from $[d]$. Each user submits an incomplete report that contains only the sampled dimensions (with added noise) in their local data. This is an effective strategy for statistical estimation, but will totally ruin the utility of a single report in PIC as unsampled dimensions are missing. Clearly, they are unsuitable for use in the PIC model.
There have been LDP randomizers that submit complete reports. One obvious strategy is to explicitly split the local budget into $d$ parts and then apply a one-dimensional LDP mechanism independently to each dimension, or implicitly distribute the budget among dimensions, as seen in the Laplace [27], PlanarLaplace [4], PrivUnit [18], and PrivUnitG mechanisms [8]. However, these approaches are sub-optimal in the high budget regime. Specifically, even if we use the optimal one-dimension randomizer [9], splitting the budget across each dimension and then applying any randomizer for each dimension will result in a mean squared error (MSE) of at least $\frac{d}{(e^{\epsilon/d}-1)^{2/3}}$. The Laplace/PlanarLaplace mechanisms introduce an MSE rate of $\frac{d}{\epsilon^{2}}$, while the PrivUnit/PrivUnitG mechanisms incur an MSE rate of $\frac{d}{\min\{\epsilon,\epsilon^{2}\}}$. In contrast, later we will show that the MSE rate can be improved to $(e^{\epsilon}-1)^{-2/(d+2)}$ (see Section VI-C and Appendix -B).
Another strategy for submitting complete reports involves using additional randomization techniques such as random projection, data sketches, public randomness, or quantization. Randomizers employing this strategy [73, 33, 71, 19, 49, 6] avoid the issue of incomplete reports but introduce additional noise because of the extra randomization. While the additional noise is not significant in the low budget regime (i.e., $\epsilon=O(1)$) that is typical in the LDP model, it becomes dominant in the PIC model where the local budget can be as large as $\tilde{O}(\log n^{\prime}_{i^{*}})$. The resulting additional error will never diminish even when $\epsilon\to+\infty$.
VI-B Randomizer Design
We now introduce an asymptotic optimal randomizer, tailored for the PIC model. The randomizer uses an LDP mechanism, which we termed as Minkowski Response. For ease of presentation, here we will focus primarily on the $\ell_{2}$ case (and the $\ell_{+\infty}$ case in Appendix -B), but the mechanism can be generalized to other Minkowski distances.
Without loss of generality, we assume the user data domain to be a $\ell_{2}$-bounded hyperball $\mathbb{X}=\{x\ |\ x\in\mathbb{R}^{d}\ and\ \|x\|_{2}\leq 1\}$. Most, if not all, real-world data domains can be normalized to $\mathbb{X}$ (e.g. gradient vector, set-valued data, and location data). We also denote an $\ell_{2}$-bounded hyperball with radius $r$ centered at any $x\in\mathbb{R}^{d}$ as follows:
$\mathbb{B}_{r}(x)=\{x^{\prime}\ |\ x^{\prime}\in\mathbb{R}^{d}\ and\ \|x^{\prime}-x\|_{2}\leq r\},$
and it is shorted as $\mathbb{B}_{r}$ when $x=\vec{0}$.
Minkowski Response works by first defining a distance $r$ based on the local privacy budget as the following:
$r=\big{(}(e^{\epsilon}-1)^{1/(d+2)}-1\big{)}^{-1}.$
Then given the input domain $\mathbb{X}$, the output domain $\mathbb{Y}_{r}$ is defined by expanding $\mathbb{X}$ by $r$:
$\mathbb{Y}_{r}=\{y\ |\ y\in\mathbb{R}^{d}\ and\ \exists x\in\mathbb{X}\ that\ y\in\mathbb{B}_{r}(x)\}.$
Figure 4: Illustrative probability design of Minkowski response mechanism over two dimensions with cap area radius $r$, given different inputs $x_{i}$ in (a) and (b) respectively.
For any input $x\in\mathbb{X}$, Minkowski Response outputs an output $y\in\mathbb{Y}_{r}$ with relatively high probability in the cap area $\mathbb{B}_{r}(x)$ and relatively low probability in remaining domain $\mathbb{Y}_{r}\backslash\mathbb{B}_{r}(x)$ (see Figure 4). Formally:
$y=\left\{\begin{array}[]{@{}ll@{}}\text{uniform}(\mathbb{B}_{r}(x)),&\text{with prob. $\frac{\mathrm{V}(\mathbb{B}_{r})\cdot(e^{\epsilon}-1)}{\mathrm{V}(\mathbb{Y}_{r})+\mathrm{V}(\mathbb{B}_{r})\cdot(e^{\epsilon}-1)}$;}\\ \text{uniform}(\mathbb{Y}_{r}),&\text{with prob. $\frac{\mathrm{V}(\mathbb{Y}_{r})}{\mathrm{V}(\mathbb{Y}_{r})+\mathrm{V}(\mathbb{B}_{r})\cdot(e^{\epsilon}-1)}$,}\\ \end{array}\right.$ (2)
where $V(*)$ denote the volume of the corresponding domain.
Lastly, $y$ is debiased to $\tilde{x}$ so that $\mathbb{E}[\tilde{x}]={x}$ as follows:
$\tilde{x}=y\cdot\frac{\mathrm{V}(\mathbb{Y}_{r})+\mathrm{V}(\mathbb{B}_{r})\cdot(e^{\epsilon}-1)}{V(\mathbb{B}_{r})\cdot(e^{\epsilon}-1)}.$ (3)
It is obvious that the output of Minkowski Response is informative in every dimension. Therefore it avoids problems brought up by incomplete reports. Also, intuitively it has a better utility because the mechanism is more likely to output a value near the true value (i.e., in $\mathbb{B}_{r}(x)$), than from other parts of the output domain (i.e. $\mathbb{Y}_{r}$). When the budget $\epsilon$ gets large, the error rate of Minkowski Response decays faster than in previous LDP mechanisms. More specifically, the decay rate of Minkowski Response is $(e^{\epsilon}-1)^{-2/(d+2)}$ (Equation 5 in Appendix -B), while that of the previous mechanisms is $d/(e^{\epsilon/d}-1)^{2/3}$ or $d/\epsilon^{2}$. Therefore, the utility advantage of Minkowski Response becomes more significant when $\epsilon$ gets larger. When $n^{\prime}_{i^{*}}\to+\infty$ (and thus $\epsilon\to+\infty$), $r$ becomes 0, and the error goes to zero (i.e. no additional error).
VI-C Analysis of Minkowski Response
The local privacy guarantee of the randomizer is presented in Theorem VI.1.
Theorem VI.1 (Local Privacy Guarantee)
Given input domain $\mathbb{X}=\mathbb{B}_{1}$, the Minkowski response mechanism defined in Equation 2 satisfies $\epsilon$-LDP.
Proof:
It is observed that the output probability distribution in Equation 2 is valid for any input $x\in\mathbb{X}$, the probability density in the cap area $\mathbb{B}_{r}(x)$ is $\frac{e^{\epsilon}}{\mathrm{V}(\mathbb{Y}_{r})+\mathrm{V}(\mathbb{B}_{r})\cdot(e^{\epsilon}-1)}$, and the density in the non-cap area $\mathbb{Y}_{r}\backslash\mathbb{B}_{r}(x)$ is $\frac{1}{\mathrm{V}(\mathbb{Y}_{r})+\mathrm{V}(\mathbb{B}_{r})\cdot(e^{\epsilon}-1)}$. Therefore, for any $x,x^{\prime}\in\mathbb{X}$, we have $\frac{\mathbb{P}[\mathcal{R}(x)=y]}{\mathbb{P}[\mathcal{R}(x^{\prime})=y]}\leq e^{\epsilon}$ for all possible $y\in\mathbb{Y}_{r}$, establishing the local $\epsilon$-DP guarantee of the Minkowski response mechanism $\mathcal{R}$. ∎
Next, we analyze the utility of the Minkowski Response in the PIC model. As previously mentioned, in the PIC model, the single report error is a more appropriate measure of utility compared to the statistical errors used in the conventional shuffle model. In Theorem VI.2, we examine the single report error in the PIC model and establish its lower bound. The proof is provided in Appendix -C.
Theorem VI.2 (Error Lower Bounds)
Given $d\in\mathbb{N}$, $\epsilon_{c}>0$, $\delta\in(0,0.5]$, $\mathbb{X}=\mathbb{B}_{1}(\{0\}^{d})$, then for any randomizer $\mathcal{R}:\mathbb{X}\mapsto\mathbb{R}^{d^{\prime}}$ such that $\mathcal{S}\circ\mathcal{R}(X)$ and $\mathcal{S}\circ\mathcal{R}(X^{\prime})$ are $(\epsilon_{c},\delta)$-indistinguishable for all possible neighboring datasets $X,X^{\prime}\in\mathbb{X}^{n}$, and for any estimator $f:\mathbb{R}^{d^{\prime}}\mapsto\mathbb{X}$, we have $\max_{x\in\mathbb{X}}\mathbb{E}\big{[}\|f\circ\mathcal{R}(x)-x\|_{2}^{2}\big{]}\geq\tilde{\Omega}\big{(}{1}/{n^{\frac{2}{d+2}}}\big{)}.$
The above theorem suggests that in the PIC model, for any randomizer, the single report error is at least $\tilde{\Omega}\big{(}{1}/{n^{\frac{2}{d+2}}}\big{)}$. Clearly, a randomizer offers better utility if its error is closer to this bound. For the randomizer using Minkowski Response, as presented in section VI-B, we can demonstrate that its single report error has an upper bound. This is summarized in Theorem VI.3, with the proof provided in Appendix -B.
Theorem VI.3 (Error Upper Bounds)
For any $d\in\mathbb{N}$, $\epsilon_{c}>0$, $\delta\in(0,0.5]$, $\mathbb{X}=\mathbb{B}_{1}(\{0\}^{d})$, if $\epsilon_{c}\leq O(1)$ and $n>\max\{16\log(1/\delta),\frac{2^{d+7}\log(1/\delta)}{(e^{\epsilon_{c}}-1)^{2}}\}$, then there exist a randomizer $\mathcal{R}:\mathbb{X}\mapsto\mathbb{R}^{d}$ such that $\mathcal{S}\circ\mathcal{R}(X)$ and $\mathcal{S}\circ\mathcal{R}(X^{\prime})$ are $(\epsilon_{c},\delta)$-indistinguishable for all possible neighboring datasets $X,X^{\prime}\in\mathbb{X}^{n}$, and:
$\max_{x\in\mathbb{X}}\mathbb{E}\big{[}\|\mathcal{R}(x)-x\|_{2}^{2}\big{]}\leq O\Bigg{(}\Big{(}\frac{\log(1/\delta)}{n\epsilon_{c}^{2}}\Big{)}^{\frac{2}{d+2}}\Bigg{)}.$
Given Theorems VI.2 and VI.3, we observe that as $n\rightarrow\infty$, the error upper bound $\tilde{O}(\frac{1}{n^{2/(d+2)}})$ of the Minkowski Response randomizer is the same as the error lower bound of the PIC model (the $\tilde{O}$ notation suppresses factors about $\epsilon_{c},\delta$ for simplicity). This means that the utility of the Minkowski Response is asymptotically optimal. In contrast, existing LDP randomizers with privacy amplification in the PIC model will suffer the single report error of $\tilde{O}(\frac{d}{n^{2/(3d)}})$ or $\tilde{O}(\frac{d}{\log^{2}n})$, whose denominator grows much slower than $n^{2/(d+2)}$ when $d>1$ and $n$ is relatively large. In practice, $n$ does not need to be very large; in our experiments, Minkowski Response outperforms existing LDP randomizers in PIC model once $n$ reaches the order of $10^{2}$. When comparing the PIC model to local DP model, the gap is undoubtedly larger, as numerical methods of privacy amplification [56, 85] will always give a local budget $\epsilon$ that is larger than the global privacy $\epsilon_{c}$ as long as $n\geq 2$, creating an extra gap in between the local and PIC models of these LDP randomizers. Actually, any randomizers in the LDP model must endures an error of $\Omega({d}/{\epsilon_{c}^{2}})$ for single report when $\epsilon_{c}\leq O(1)$ [18, 25].
VII Experimental Evaluation
We evaluate the utility and efficiency performances of proposed PIC model in three representative individual computation tasks: spatial crowdsourcing, location-based nearest neighbors, and federated learning with incentives, aim to answer the following research questions:
  • RQ1.
    How much do our proposals outperform existing approaches in accuracy/utility?
  • RQ2.
    What the computational/communication costs of our model for supporting security functionalities?
TABLE II: Mean $\ell_{2}$-error (and running time) comparison of local $\epsilon$-LDP randomizers on location domain $[-1,1]\times[-1,1]$. All results are the expected value of $1000$ repeated experiments.
randomizer $\epsilon=0.5$ $\epsilon=1.0$ $\epsilon=2.0$ $\epsilon=3.0$ $\epsilon=5.0$ $\epsilon=8.0$ $\epsilon=10.0$
PrivUnit [16] 8.94 (49us) 4.68 (42us) 2.25 (61us) 1.44 (63us) 0.81 (113us) 0.32 (305us) 0.18 (868us)
PrivUnitG[8] 8.73 (76ms) 4.63 (75ms) 2.27 (77ms) 1.51 (76ms) 0.96 (74ms) 0.63 (75ms) 0.53 (81ms)
Laplace [28] 12.97 (0.1us) 6.56 (0.1us) 3.27 (0.1us) 2.13 (0.1us) 1.30 (0.1us) 0.81 (0.1us) 0.64 (0.1us)
PlanarLaplace [4] 11.17 (12us) 5.63 (11us) 2.84 (11us) 1.88 (11us) 1.14 (11us) 0.71 (12us) 0.56 (12us)
Staircase[35] 13.19 (49us) 6.40 (48us) 3.13 (46us) 2.01 (48us) 1.05 (44us) 0.46 (51us) 0.28 (50us)
SquareWave[57] 11.87 (1.9us) 5.72 (2.5us) 2.65 (1.9us) 1.68 (2.2us) 0.92 (1.7us) 0.53 (2.2us) 0.42 (2.1us)
MinkowskiResponse 10.42 (0.7us) 4.50 (0.6us) 1.78 (0.8us) 0.98 (0.8us) 0.39 (0.6us) 0.14 (0.7us) 0.074 (0.8us)
VII-A Spatial Crowdsourcing
Datasets and settings. In the experiments for spatial crowdsourcing, we use use two real datasets, GMission dataset [20] and EverySender dataset [79], for scientific simulation and campus-based micro-task completion respectively. The number of users/workers, location domain range, and serving radius of workers about these two datasets are summarized in Table III (more information are presented in Appendix -F). In our experiments, for compliance with data randomizers (see following parts), we rescale the location domain range to $[-1,1]\times[-1,1]$ and rescale the corresponding serving radius to $0.41.0\cdot\frac{1-(-1)}{5-0}=$ 0.4.
TABLE III: Summary statistics of spatial crowdsourcing datasets.
Dataset users workers location domain serving radius
GMission 713 532 $[0,5.0]\times[0,5.0]$ $1.0$
EverySender 4036 817 $[0,5.0]\times[0,5.0]$ $1.0$
Cryptography implementation. Follow the protocol in Section V, we implement the PIC model using Elliptical Curve Diffie-Hellman key agreement (ECDHE [11]) for establishing shared secret between client/server and matched user/worker, meanwhile using AES encryption [68] and HMAC signature [14] for secure point-to-point communication after key agreement.
Data randomizers and server-side algorithms. We use the Minkowski response for location randomization. As comparison, in the local model of DP (e.g., in [83, 77, 81]), we compare with existing mechanisms including Laplace [28], Staircase [35], PlanarLaplace with geo-indistinguishability [4], SquareWave [57], PrivUnit [16] and its Gaussian variant PrivUnitG in [8]. On the server side, we use either minimum weighted full matching [51] or maximum matching [45] over bipartite graph as the user/worker matching algorithm based on (noisy) locations, to obtain a match $M:G_{a}\times G_{b}\mapsto\mathbb{R}$. Specifically, the minimum weighted full matching aims to minimize the overall travelling costs between users and works over the (noisy) fully-connected bipartite graph; the maximum matching aims to maximize the number of successfully matched user/worker pairs over the (noisy) partially-connected bipartite graph, where user/worker pairs with distance surpassing $0.4$ (i.e. the serving radius) are deemed as not connected.
Performance metrics. We evaluate the performance of spatial crowdsourcing in terms of travel costs, successful matching ratio, and running time. The travel costs are the total actual Euclidean distance between all pairing matched user/workers:
$\text{travel costs}=\sum\nolimits_{(i,j)\in G_{a}\times G_{b}}\llbracket M(i,j)>0\rrbracket\cdot\|l_{i}-l_{j}\|_{2},$
[…]
$\text{success ratio}=\frac{\sum_{(i,j)\in G_{a}\times G_{b}}\llbracket M(i,j)>0\rrbracket\cdot\llbracket\|l_{i}-l_{j}\|_{2}\leq\tau\rrbracket}{\min\{|G_{a}|,|G_{b}|\}}.$
We evaluate both the client-side and server-side running time of our protocol on a laptop computer embedded with Intel i5555-8250U CPU @1.6GHz and $8$GB memory.
Figure 5: Expected $\ell_{2}$ error of reported locations.
Performance of location reports. Before exploring spatial crowdsourcing tasks, we first evaluate our proposed Minkowski response mechanism against existing mechanisms, as presented in Table II. This evaluation focuses on errors and running times associated with reporting locations under local $\epsilon$-DP constraints. We quantify the error using the expected $\ell_{2}$ distance between each true and noisy location pair. Notably, the Minkowski response showcases an $\ell_{2}$ error comparable to state-of-the-art (SOTA) mechanisms when $\epsilon\leq 1.0$. Its superiority is evident when $\epsilon\geq 2.0$, displaying a significantly reduced error. This supports the theoretical assertions made in Section VI, underscoring the importance of the Minkowski response for the shuffle model, over other existing randomizers. Additionally, the Minkowski response is highly efficient, demanding less than $1$us on the user end. This speed is nearly equivalent to merely adding Laplace noises and is roughly $50$x faster than the previously SOTA PrivUnit.
We further contrast the utility performances of the mechanisms in the local model with those in the shuffle model. Privacy amplification in the shuffle model is achieved using the latest analyses from [32]. As illustrated in Figure 5, the shuffle model presents a significantly reduced location reporting error compared to the local model. This disparity becomes more pronounced as the number of users increases (notably, the EverySender dataset comprises more users than GMission). The introduction of an apt randomizer—such as the Minkowski response—can further elevate utility. This is evident when comparing the Minkowski response and PrivUnit, both of which operate within the shuffle model.
Performance of minimum weight matching. Figure 6 presents the expected travel costs associated with minimum weighted full matching. For a comprehensive overview of the experimental results involving more competitors, readers are directed to -F. Notably, the Squarewave mechanism outperforms PrivUnit in matchings, prompting its primary representation in the shuffle model discussions. The equivalent shuffle model yields a cost reduction of approximately $50\%$, with the savings escalating as the number of users increases. The Minkowski response consistently exhibits the lowest costs across all scenarios. Together, they result in a substantial reduction of about $75\%$ when compared to existing approaches.
Figure 6: Travel costs of minimum weighted matching.
Performance of maximum matching. Figure 7 illustrates the success ratios for maximum matching. The equivalent shuffle model shows a relative improvement of approximately $200\%$ in the success ratio. The Minkowski response consistently yields higher ratios across all scenarios. When the user count becomes relatively substantial (for instance, a few thousands as in EverySender), our results indicate that spatial crowdsourcing maximum matching, with a reasonable level of privacy protection (specifically, $\epsilon\leq 3$), can be made practical, achieving a success ratio exceeding $90\%$.
Figure 7: Success ratios of maximum matching.
TABLE IV: Running time and communication overheads of spatial crowdsourcing in the PIC model.
User-side Procedures Time Communication Server-side Procedures Time Communication
location randomization 0.7us - decrypt one message 2.1ms -
encrypt information 2.9ms - min-weight match (GMission) 31ms 1.2MB
send message to shuffler - 1.3KB maximum match (GMission) 15ms 1.2MB
retrieve matching result - 1.8KB min-weight match (EverySender) 79ms 4.3MB
key agreement with worker 2.6ms - maximum match (EverySender) 63ms 4.3MB
Running time and communication overheads. We present the computation and communication overheads of the PIC model in Table IV. When considering multiple procedures on the user side, the total running time for each user is only a few milliseconds (with the data randomization time being negligible), and the total communication overhead is under $4$KB. On the server side, the running time of the matching algorithm on noisy plaintext is consistent with non-private settings, scaling with the number of users/workers and taking dozens of milliseconds. The additional decryption/encryption runtime is capped at several seconds, mirroring the costs of HTTPS decryption/encryption. Specifically, it utilizes ECDHE for key agreement and AES encryption as with HTTPS on TLS 1.3 [67] for client-server communication. In essence, the running times on both the client and server sides of the PIC model closely resemble those in non-private settings that use HTTPS communication (for a fair comparison).
VII-B Location-based Social Systems
In this part, we evaluate the performance of PIC model for privacy preserving nearest neighbor (NN) queries on users, a prevalent task in location-based social networks [22]. Like the exemplar spatial crowdsourcing tasks, the nearest location senders’ information (including their corresponding public keys) w.r.t. $\ell_{2}$-distance) of a user is securely return to the user. These matched neighboring users can then securely communicate with other to accomplish subsequent tasks.
TABLE V: Statistics of location social network datasets.
Dataset users check-ins location domain
Gowalla 138368 6442892 $[37.54,37.79]\times[-122.51,-122.38]$
Foursquare 1083 227428 $[40.55,40.99]\times[-74.27,-73.68]$
Datasets and settings. We use use two real check-in datasets, Gowalla dataset [22] and Foursquare dataset [87]. The statistics of two datasets are summarized in Table V. The location domains are rescaled to the standard domain $[-1,1]\times[-1,1]$. As for nearest neighbor queries, we set the nearest radius $\tau$ for Gowalla to $0.1$ and the nearest radius $\tau$ for Foursquare to $0.2$, so that users have a few tens or hundreds neighbors (vary due to check-ins densities). Accordingly, the population size for privacy amplification in PIC model is set to $n-1000$. The concrete cryptography implementations and data randomizers for nearest neighbor queries are the same as the one for spatial crowdsourcing, while the server simply runs a radius-based nearest neighbor algorithm and return the neighboring (public key, noisy location) lists to each user.
Performance metrics. We evaluate the performance of nearest neighboring queries in terms of precision and recall, and $F_{1}$ score of neighboring lists. Let $N_{i}$ denote the true neighbor lists that have $\ell_{2}$ distance to user $i$ less than the parameter $\tau$, and let $\widehat{N}_{i}$ denote the retrieved neighbor lists that have $\ell_{2}$ distance to user $i$ based on noisy (sanitized) locations less than the parameter $\tau$. The precision of nearest neighboring queries is defined as $\text{precision}={(\sum_{i\in[n]}|N_{i}\bigcap\widehat{N}_{i}|)}/{(\sum_{i\in[n]}|N_{i}|)}$, the recall of nearest neighboring queries is $\text{recall}={(\sum_{i\in[n]}|N_{i}\bigcap\widehat{N}_{i}|)}/{(\sum_{i\in[n]}|\widehat{N}_{i}|)}$. We use a balanced indicate
$F_{1}\text{ score}=2\cdot\text{precision}\cdot\text{recall}/(\text{precision}+\text{recall})$
as the major measurement. All presented results are the average value of $10$ experimental runs.
Experimental results. Figure 8 displays the $F_{1}$ scores for radius nearest neighbor queries. To evaluate the effects of varying user numbers, we adjusted the number of users $n$ in the Gowalla dataset from $10000$ to $40000$. Notably, as $n$ surpasses $50000$, the PIC model’s scores see an approximate relative increase of $100\%$ compared to the local model. The Minkowski response consistently records the highest scores. When used in conjunction, they can achieve up to $90\%$ of the $F_{1}$ scores within a reasonable privacy loss (e.g., $\epsilon\leq 3$), especially when $n$ is relatively large.
Figure 8: F1 scores of radius nearest neighbor queries in location-based social systems.
VII-C Federated Learning with Incentives
Recall that in federated learning with rigorous privacy protection, many clients collaboratively contribute gradient updates of model parameters so as to build a machine learning model. Each client $i$ holds a local dataset $D_{i}$. In each iteration $t$, a client $i$ downloads the latest model parameter $W_{t-1}$, and computes the local gradient $g_{i}$ based on local dataset $D_{i}$. Then, the client sanitizes $g_{i}$ and sends it to the shuffler/server. Without loss of generality, we assume the server uses Shapley value [69] to determine the gradient payoffs, where the utility function for computing Shapley value is based on the model parameters’ performance (e.g., accuracy or the negative of loss function) on a validation dataset $D_{val}$. The Shapley values themselves may severely leak sensitive information about user data [59]. Despite the vast non-linear computations involved in evaluating Shapley values with neural networks, no rigorous protection has been offered for federated learning with incentives in existing literature. In this study, we demonstrate the feasibility of computing Shapley value without incurring additional privacy loss.
TABLE VI: Model architecture with 4292 trainable parameters.
Layer Parameters
Convolution 8 filters of 4$\times$4, stride 2
Max-pooling 2$\times$2
Convolution 6 filters of 5$\times$5, stride 2
Max-pooling 2$\times$2
Softmax 10 units
Datasets and settings. Our experiments with federated learning use the MNIST handwritten digit dataset, comprising $60,000$ training, $5,000$ validation, and $5,000$ test images. We employ a neural network outlined in Table VI with $d=4292$ trainable parameters. It achieves $98.6\%$ accuracy using non-private, untruncated vanilla SGD and $92.7\%$ accuracy with non-private, truncated SGD. For our experiments, we simulate $50,000$ clients with each client possessing one sample (i.e., $n=50,000$). During each private SGD step, $s=10,000$ clients are randomly selected. Each client subsamples $0.5\%$ of the gradient vector dimensions, then clips the $\ell_{1}$-norm of the block gradient with a threshold of $c=0.001$. We compare our PIC model against a local model, both implementing Laplace mechanism for gradient randomization.
Implementation details. We use Elliptical curve encryption to perform secure point-to-point information delivery. Specifically, to delivering monetary incentives with anonymity protection, clients and server may employ untraceble crypto-currency (e.g., ZCash [46]). Each client creates one random temporary wallet address and append it to the message, the server computes incentive payments of the message and then send corresponding amount of coins to the wallet address.
Performance metrics. We assess performances using gradient/model accuracy and Shapley value accuracy. Gradient accuracy is gauged using the mean squared error (MSE) between the subsampled, clipped, and privatized gradients $\hat{g}_{i}$ and the corresponding non-private gradients ${g}_{i}$:
$\text{gradient}\ MSE=\|\sum\nolimits_{i\in[s]}\hat{g}_{i}-\sum\nolimits_{i\in[s]}{g}_{i}\|_{2}^{2}.$
Model accuracy refers to the image classification accuracy on the test dataset. Shapley value accuracy is the MSE between Shapley values of privatized gradients $\hat{g}_{i}$ and non-private gradients ${g}_{i}$:
$\text{shapley}\ MSE=\sum\nolimits_{i\in[s]}(\widehat{\texttt{Shapley}}_{i}-{\texttt{Shapley}}_{i})^{2},$
where the Shapley values for privatized gradients use the utility function (see Appendix -H for detail):
$\widehat{U}(S)=\frac{<\texttt{grad}_{val},\sum_{i\in S}\hat{g}_{i}>}{\|\texttt{grad}_{val}\|_{2}\cdot\|\sum_{i\in S}\hat{g}_{i}\|_{2}}.$
The presented gradient/shapley MSE averages over $50$ aggregation rounds, and the model accuracy is from the final round.
Experimental results. Since the curator model of DP and other cyrptographic MPC approaches (e.g., secure aggregation [18]) are not applicable to privacy preserving joint gradient aggregation and Shapley value computation (with complicate non-linear computation), we compare equivalent shuffle model with the local model of DP. Both models use Laplace noises [28] to privatizing local gradients, and use tight numerical composition for sequential gradient averaging [42]. We present the experimental results in Table VII, with varying overall privacy budget $\epsilon$. Every result is the average value of $5$ simulations.
TABLE VII: Utility performance of joint gradient aggregation and Shapley-value incentive for federated learning with differential privacy guarantee on MNIST dataset.
Setting
Privacy
Budget
Test
Accuracy
Gradient
MSE
Shapley
MSE
local model $\epsilon=1$ $19.1\%$ $1.7$e-4 $8.9$e-3
$\epsilon=3$ $28.9\%$ $1.3$e-4 $6.1$e-3
$\epsilon=5$ $35.6\%$ $1.1$e-4 $5.2$e-3
PIC model $\epsilon=1$ $77.6\%$ $2.5$e-5 $4.6$e-3
$\epsilon=3$ $81.2\%$ $1.4$e-5 $4.2$e-3
$\epsilon=5$ $85.1\%$ $1.1$e-5 $3.9$e-3
It is observed that PIC model enjoys significantly better gradient accuracy, increase model test accuracy by about $50\%$ in the absolute sense, reduces about $90\%$ gradient aggregation error, when compared to the local model. The PIC model also reduces about $40\%$ error on Shapley values, permit accurate payoff distribution to incentive honest/informative participants.
VIII Conclusion
In this paper, we proposed private individual computation model, for multi-party individual computation with DP preservation, having broad application to spatial crowdsourcing, resource allocation, location-based systems, and federated learning with incentives. The paradigm introduces statistically random identities into the traditional shuffle model of DP, so as to support secure functionalities required by individual computations, meanwhile maintaining the most benefits of privacy amplification via shuffling. It incurs only $O(d)$ computation and communication costs (matching the non-private setting) on the client side, meanwhile permitting plaintext computation on the server side. Additionally, with the increasing of number of participating parties, the utility of individual computation grows (as there are stronger privacy amplification effects), making it suitable for large-scale applications.
The simple paradigm can efficiently solves prevalent tasks that are computationally intractable in MPC approaches (e.g., Ishai’s anonymous model, secure aggregation, or garbled circuits), such as federated deep learning with incentives and large-scale spatial crowdsourcing, necessitating essentially the same overheads as in non-private settings. Moreover, it reduces $40\%$-$90\%$ errors/costs, increases $100\%$-$300\%$ utility, when compared to existing tractable approaches (e.g., local DP).
Discussions and future research. The PIC model guarantees privacy gracefully in cases of malicious clients or collusion, and can adapt to fully decentralized settings without an orchestrating server. Given its alignment with random digital identities in Blockchains and in Ad tracking systems [54], it offers potential to enhance formal privacy in these systems.