### v2

# Beyond Statistical Estimation: Differentially Private Individual Computation via Shuffling

Abstract
In data-driven applications, preserving user privacy while enabling valuable computations remains a critical challenge. Technologies like Differential Privacy (DP) have been pivotal in addressing these concerns. The shuffle model of DP requires no trusted curators and can achieve high utility by leveraging the privacy amplification effect yielded from shuffling. These benefits have led to significant interest in the shuffle model. However, the computation tasks in the shuffle model are limited to statistical estimation, making the shuffle model inapplicable to real-world scenarios in which each user requires a personalized output. This paper introduces a novel paradigm termed Private Individual Computation (PIC), expanding the shuffle model to support a broader range of permutation-equivariant computations. PIC enables personalized outputs while preserving privacy, and enjoys privacy amplification through shuffling. We propose a concrete protocol that realizes PIC. By using one-time public keys, our protocol enables users to receive their outputs without compromising anonymity, which is essential for privacy amplification. Additionally, we present an optimal randomizer, the Minkowski Response, designed for the PIC model to enhance utility. We formally prove the security and privacy properties of the PIC protocol. Theoretical analysis and empirical evaluations demonstrate PIC’s capability in handling non-statistical computation tasks, and the efficacy of PIC and the Minkowski randomizer in achieving superior utility compared to existing solutions.
Index Terms:
Differential Privacy, Private Individual Computation, Shuffle Model, Privacy-Utility Trade-offs
I Introduction
Personal information fuels a wide array of data-driven applications, e.g. statistical analytics, machine learning, recommendation systems, spatial crowdsourcing, e-health, social networks, and smart cities. These applications deliver substantial value but rely on data collected from users, which is a prime target for attacks and carries a high risk of leakage or abuse. Data privacy concerns are escalating, especially after several high-profile data breach incidents. Despite the introduction of stricter privacy laws such as the EU’s General Data Protection Regulation and the California Consumer Privacy Act, many users still distrust service providers and are hesitant to consent to the use of their data. To bridge this trust gap and encourage user participation, significant efforts have been made recently to develop privacy-enhancing technologies that enable private data processing.
Two prominent technologies addressing this problem are secure multiparty computation (MPC) [91] and differential privacy (DP) [31]. MPC employs interactive cryptographic protocols that enable mutually untrusted parties to jointly compute a function on their private data, ensuring each party receives an output while learning nothing else. Despite its strong privacy guarantees, the cryptographic nature of MPC results in substantial overhead, posing scalability challenges for practical applications. In contrast, DP enhances privacy by adding random noise to the data, allowing computation to be performed on sanitized data without the need for heavy cryptographic machinery. Traditional central DP relies on a trusted curator [31], but the local model (LDP [54]) enables users to sanitize their data locally before sharing it. LDP’s practicality and minimal trust requirements have led to its adoption by companies such as Apple [77], Microsoft [27], and Google [33] in their real-world systems. However, LDP’s primary drawback is that each user must independently add sufficient random noise to ensure privacy, which can significantly impact the utility of the data.
Recently, within the realm of DP research, the shuffle model [13, 32] has emerged. This model introduces a shuffler that randomly permutes messages from users and then sends these anonymized messages to a computing server or analyzer. The server can then compute on these messages to derive the result. Trust in the shuffler is minimized because each user encrypts their locally sanitized data using the server’s public key. This way, the shuffler remains oblivious to the messages received from the users. A key advantage of the shuffle model over LDP is utility. It has been proven that anonymizing/shuffling messages amplifies the privacy guarantee provided by the local randomizer used by the users. For instance, shuffled messages from $n$ users each adopting local $\epsilon$-DP actually preserve differential privacy at the level $\epsilon_{c}=\tilde{O}(\sqrt{e^{\epsilon}/n})$ [34, 35]. Consequently, to achieve a predefined global privacy goal, less noise must be added when users sanitize their data locally. This significantly improves the accuracy of the final result. Owing to these utility advantages, extensive studies have been conducted within the shuffle model, e.g., [8, 40, 41, 41, 42].
That said, the shuffle model has a noticeable limitation. The privacy amplification effect relies heavily on anonymizing/shuffling messages, which significantly restricts the types of computation that can be performed. So far, the sole form of computation achievable within the shuffle model is statistical estimation, i.e., the server takes the shuffled messages, aggregates them, and computes a single output from them, e.g., a count, sum, or histogram. However, many real-world applications are non-statistical in nature. When multiple users pool their data together for joint computation, they expect an individualized output that may differ for each user. We coined the term “individual computation” for such tasks. Examples of individual computation tasks include:
  • I.
    Combinatorial optimization: Spatial crowdsourcing [83], advertisement allocation [62], and general combinatorial optimization [56], where two or more parties are often matched together based on their private information. Each party should get their own list of “best matches” whatever that means.
  • II.
    Information retrieval: Mobile search [53], location-based systems [2, 21], where the query results (e.g., nearby restaurants or neighboring users) depend on the private information of the inquirer.
  • III.
    Incentive mechanisms: In federated learning [92] or data crowdsensing [74], incentives play a vital role to encourage well-behaved participation. The amount of rewarding incentive must be computed for individual users based on their contribution (e.g., via Shapley values [71]).
At first glance, private individual computation appears unattainable within the shuffle model because the need for personalized output conflicts with the anonymization required for privacy amplification. However, this is not necessarily the case. Our observation is that many individual computing tasks are equivariant to shuffling, meaning that the permutation applied to the inputs does not affect the computation. Therefore, shuffling does not prevent the server from producing personalized answers for each client – the server does not need to know which answer belongs to whom. Nevertheless, there is a challenge: how to return the output to the correct user without compromising anonymity. One straightforward approach is to have the shuffler maintain a long-term duplex connection channel between each client and the server (e.g., as in an Onion routing network [44]). However, this method is costly due to the need to store communication states and may be vulnerable to de-anonymization attacks on anonymous channels [64, 66]. Additionally, current shuffler implementations, e.g. the one described in the seminal work [13], do not support such duplex connections. Moreover, after receiving the computation results, many individual computation tasks require establishing party-to-party communication (e.g., a user communicates with the matched driver in taxi-hailing services, a user communicates with matched near users in social systems), which is intractable in the duplex shuffle channel. Addressing this issue is the first technical challenge we need to overcome.
The second challenge we face is designing optimal randomizers for individual computation within the shuffle model. Often, the randomizer in a shuffle model protocol should be tailored to specific tasks. For statistical tasks within the shuffle model, several studies have developed near-optimal randomizers, as demonstrated in histogram estimation [39, 34] and one-dimensional summation estimation [7]. However, the new setting of individual computation is different: the focus is on the accuracy of the output for each user rather than the statistical accuracy of the population. This difference renders existing randomization strategies (e.g., randomizers utilizing dimension sampling, budget splitting, or data sketching, as reviewed in [88]), as well as prevalent randomizers (such as adding Laplace [31] or Staircase [38] noises), less effective for the new setting. Therefore, we need to reconsider the fundamental privacy-utility trade-offs and redesign the underlying randomizers to better suit individual computation requirements.
Our contributions In this work, we introduce a new paradigm extending the shuffle model that allows a wider range of permutation-equivariant tasks to be computed with DP guarantee and can enjoy privacy amplification provided by shuffling. We term the new paradigm Private Individual Computation (PIC in short). We define PIC formally as an ideal functionality, which captures its functional and security properties. We also provide a concrete protocol, with formal proof, that can realize the ideal functionality.
Similar to the shuffle-DP protocols, each user adds noise locally to their data and encapsulates it (and possibly other auxiliary data) into an encrypted message under the public key of the computation server. The shuffler then shuffles the encrypted messages, before sending them to the computation server. The server can decrypt these messages and perform the permutation-equivariant computation. What differs is that each user also includes a one-time public key within the encrypted message. This one-time public key serves two purposes: (1) it allows the server to encrypt the computation result such that it can only be decrypted by the owner of the corresponding private key, and (2) it acts as a pseudonym for the key owner. This approach addresses our first challenge with minimal overhead: the server can publish a list where each entry consists of a public key along with the computation result encrypted under this key. Users can download the entire list and decrypt the entry associated with their own public key, maintaining their anonymity. Additionally, users can establish secure communication channels with other matched parties using their public keys to eventually complete PIC tasks.
Another main contribution is the development of an asymptotically optimal randomizer specifically designed for the PIC model. This randomizer is based on an LDP mechanism we call Minkowski Response. The primary goal of this design is to enhance utility, measured by the single-report error, which is the expected squared error between a user’s true data value and its sanitized version. To achieve this, a Minkowski distance $r$ is determined based on the privacy budget. The randomizer’s output domain is defined as a ball extending the input domain’s radius by $r$. The key to achieving high utility lies in the randomizer’s output: it selects a value close to the true value (within $r$) with a relatively high probability and a value from elsewhere in the entire output domain with a relatively low probability.
We formally prove the security and privacy properties of the PIC protocol. Additionally, we provide a theoretical analysis of the utility bounds achievable by protocols in the PIC model and the Minkowski Response mechanism. Our analysis demonstrates that asymptotically, the error upper bound of the Minkowski Response matches the error lower bound for all possible randomizers in the PIC model, thereby achieving optimality. Alongside theoretical analysis, we conducted extensive experiments using real-world applications and datasets. The evaluation confirms that computations conducted in the PIC model exhibit significantly better utility than those in the LDP model. Furthermore, the performance of the Minkowski randomizer, measured by single-report error and task-specific utility metrics, surpasses that of existing LDP randomizers commonly used in the shuffle model.
Organization. The remainder of this paper is organized as follows. Section II reviews related works. Section III provides preliminary knowledge about privacy definitions and security primitives. Section IV formalizes the problem setting. Section V presents the PIC protocol. Section VI provides optimal user-side randomizers. Section VII evaluates the utility and efficiency performances of our proposals. Finally, Section VIII concludes the paper.
II Related Works
This section reviews various approaches to private computation, primarily concentrating on non-statistical tasks.
II-A Secure Multiparty Computation
Secure Multiparty Computation (MPC) is a fundamental cryptographic primitive that enables multiple parties to jointly compute a function over their inputs while no party learns anything beyond their own input and the final output of the computation. MPC was first conceptualized by Andrew Yao in the 1980s, and it has been proven that any computable function can be realized by MPC [90]. MPC relies on cryptographic protocols that exchange encrypted messages among parties. To allow computing on encrypted data, primitives such as homomorphic encryption [67, 19], secret sharing [14, 26], or garbled circuits [91] can be employed. The primary challenge in MPC lies in balancing privacy and efficiency. While MPC offers strong privacy guarantees, it often suffers from significant computational and communication overheads, which makes scaling to large datasets or numerous parties difficult. Recent research in MPC has been focusing on optimizing protocols and practical implementations [23, 55, 16, 24, 70, 25, 75, 15]. Despite a significant improvement, MPC still faces efficiency issues that hinder its widespread real-world deployment.
II-B Curator and Local DP Methods
Many works study matching, allocation, or general combinatorial optimization problems within the curator DP model [31] in the presence of a trusted party collecting raw data from clients (e.g., in [61, 22, 79, 78]). Since the assumption of a trustworthy party is often unrealistic in decentralized settings, many studies adopt the local model of DP (e.g., in [85, 68, 80, 84]), where each client sanitizes data locally and sends the noisy data to the server for executing corresponding matching/allocation algorithms. As each client must injects sufficient noises into data to satisfy local DP, the execution results often maintain low utility.
II-C Shuffle Model of DP
The recently proposed shuffle model [13, 32] combines the advantages of the curator model (e.g., high utility) and the local model (e.g., minimal trust). Depending on the number of messages each client can send to the intermediate shuffler, the shuffle model can be categorized as single-message [7, 32, 35] and multi-message [8, 39]. The single-message shuffle model leverages privacy amplification via shuffling to enhance data utility compared to the local model. A substantial body of work [32, 34, 35] demonstrates that $n$ shuffled messages from clients, each adopting a same $\epsilon$-LDP randomizer, can actually preserve $\tilde{O}(\sqrt{e^{\epsilon}/n})$-DP. By removing the constraint of sending one message, the multi-message shuffle model can achieve better utility than the single-message model and might be comparable to the curator model (e.g., in [8, 39]). However, each multi-message protocol is tailored to a specific statistical query (e.g., summation), rendering them unsuitable for permutation-equivariant tasks with non-linear computations. There is a line of works on the shuffle model for private information retrieval (e.g., in [49, 48, 37] with cryptography security and in [81, 1] with statistical DP), where the query is represented as multiple secret shares before sent to the shuffler, and the server holding the database entries returns linear-transformed entries for each query, using the duplex shuffled communication channel. This kind of duplex-communication shuffle model can be vulnerable to anonymity attacks [64, 66], and is pertained to the linear computation in private information retrieval. It can not be applied to other PIC tasks (such as combinatorial optimization and federated learning with incentive) that involve with non-linear computations, and can not provide secure user-to-user communication needed in tasks like spatial crowdsourcing and social systems.
Overall, existing works in the shuffle model primarily focus on statistical queries. This work, for the first time, explore the shuffle model for non-statistical applications (i.e., combinatorial optimization, location-based social systems, and incentive mechanisms).
II-D Combining Cryptography and DP
While cryptographic tools can protect data secrecy during multiparty computation, they do not necessarily preserve output’s privacy. DP can be employed to enhance the privacy of the outputting result of secure multiparty computation through decentralized noisy addition [45]. To account for privacy loss due to intermediate encrypted views in MPC, researchers have proposed the relaxed notion of computational DP [63] against polynomial-time adversaries. Computational DP protocols often inherited the computation/communication complexity of MPC (refer to approaches in Table IX).
III Preliminaries
In this section, we provide a concise introduction to the preliminaries. A list of notations used throughout the paper can be found in Table I.
TABLE I: List of notations.
Notation Description
$[i]$$\{1,2,...,i\}$
$[i:j]$$\{i,i+1,...,j\}$
$\mathcal{S}$ the shuffling procedure
$\mathcal{R}$ the randomization algorithm
$G_{i}$ the $i$-th group of users ($i\in[m]$)
$n_{i}$ the number of users in group $G_{i}$
$u_{i,j}$ the $j$-th user in group $G_{i}$ where $j\in[n_{i}]$
$d$ the dimension of user’s data
$\mathbb{X}$ the domain of input data
$\mathbb{Y}$ the domain of a sanitized message
$\epsilon$ the local privacy budget
$\epsilon_{c}$ the amplified privacy level
$sk$, $pk$ the secret key and the public key, respectively
$\lambda$ the security parameter of cryptography
III-A Privacy Definitions
Definition III.1 (Hockey-stick divergence [72])
For two probability distributions $P$ and $Q$, the Hockey-stick divergence between them with parameter $e^{\epsilon}$ is as follows:
$D_{\epsilon}(P||Q)=\int_{z\in\mathcal{Z}}\max\{0,P(z)-e^{\epsilon}Q(z)\}\mathrm{d}z.$
Differential privacy imposes divergence constraints on output probability distributions with respect to changes in the input. In the curator model of differential privacy, a trusted party collects raw data $x_{i}\in\mathbb{X}$ from all users to form a dataset $T=\{x_{1},\ldots,x_{n}\}$ and applies a randomization algorithm $\mathcal{R}$ to release query results $\mathcal{R}(T)$. For two datasets $T$ and $T^{\prime}$ of the same size and differing in only one element, they are referred to as neighboring datasets. Differential privacy ensures that the Hockey-stick divergence between $\mathcal{R}(T)$ and $\mathcal{R}(T^{\prime})$ is bounded by a sufficiently small value (i.e., $\delta=O(1/n)$). The formal definition of curator differential privacy is as follows:
Definition III.2 ($(\epsilon,\delta)$-DP [31])
A randomization mechanism $\mathcal{R}$ satisfies $(\epsilon,\delta)$-differential privacy iff $\mathcal{R}(T)$ and $\mathcal{R}(T^{\prime})$ are $(\epsilon,\delta)$-indistinguishable for any neighboring datasets $T,T^{\prime}\in\mathbb{X}^{n}$. That is, $\max(D_{\epsilon}(\mathcal{R}(T)\|\mathcal{R}(T^{\prime})),D_{\epsilon}(\mathcal{R}(T^{\prime})\|\mathcal{R}(T)))\leq\delta$.
In the local model of differential privacy, each user applies a randomization mechanism $\mathcal{R}$ to their own data $x_{i}$, with the objective of ensuring that the Hockey-stick divergence between $\mathcal{R}(x)$ and $\mathcal{R}(x^{\prime})$ is 00 for any $x,x^{\prime}\in\mathbb{X}$ (see Definition III.3).
Definition III.3 (local $\epsilon$-DP [54])
A randomization mechanism $\mathcal{R}$ satisfies local $\epsilon$-DP iff $D_{\epsilon}(\mathcal{R}(x)||\mathcal{R}(x^{\prime}))=0$ for any $x,x^{\prime}\in\mathbb{X}$.
The data processing inequality is considered a key feature of distance measures (e.g., Hockey-stick divergence) used for evaluating privacy. It asserts that the privacy guarantee cannot be weakened by further analysis of a mechanism’s output.
Definition III.4 (Data processing inequality)
A distance measure $D:\Delta(\mathcal{S})\times\Delta(\mathcal{S})\to[0,\infty]$ on the space of probability distributions satisfies the data processing inequality if, for all distributions $P$ and $Q$ in $\Delta(\mathcal{S})$ and for all (possibly randomized) functions $g:\mathcal{S}\to\mathcal{S^{\prime}}$,
$D(g(P)\|g(Q))\leq D(P\|Q).$
III-B The Classical Shuffle Model
Following the conventions of the randomize-then-shuffle model [20, 7], we define a single-message shuffle protocol $\mathcal{P}$ as a list of algorithms $\mathcal{P}=(\mathcal{R},\mathcal{A})$, where $\mathcal{R}:\mathbb{X}\to\mathbb{Y}$ is local randomizer on client side, and $\mathcal{A}:\mathbb{Y}^{n}\to\mathbb{Z}$ is the analyzer on the server side. We refer to $\mathbb{Y}$ as the protocol’s message space and $\mathbb{Z}$ as the output space. The overall protocol implements a mechanism $\mathcal{P}:\mathbb{X}^{n}\to\mathbb{Z}$ as follows. User $i$ holds a data record $x_{i}$ and a local randomizer $\mathcal{R}$, then computes a message $y_{i}=\mathcal{R}(x_{i})$. The messages $y_{1},\ldots,y_{n}$ are shuffled and submitted to the analyzer. We denote the random shuffling step as $\mathcal{S}(y_{1},\ldots,y_{n})$, where $\mathcal{S}:\mathbb{Y}^{n}\to\mathbb{Y}^{n}$ is a shuffler that applies a uniform-random permutation to its inputs. In summary, the output of $\mathcal{P}(x_{1},\ldots,x_{n})$ is represented by $\mathcal{A}\circ\mathcal{S}\circ\mathcal{R}(X)=\mathcal{A}(\mathcal{S}(\mathcal{R}(x_{1}),\ldots,\mathcal{R}(x_{n})))$.
In particular, when all users adopt an identical $\epsilon$-LDP mechanism $\mathcal{R}$, recent works [34, 35] have derived that $n$ shuffled $\epsilon_{l}$-LDP messages satisfy $(O((1-e^{-\epsilon})\sqrt{e^{\epsilon}\log(1/\delta)/n}),\delta)$-DP. We denote the amplified privacy level as:
$\epsilon_{c}=\texttt{Amplify}(\epsilon,\delta,n),$
the tight value of which can also be numerically computed [57, 87].
III-C Public Key Encryption
[…]
  • Gen: takes as input the security parameter $\lambda$ and outputs a pair of keys $(pk,sk)$, where $pk$ denotes the public key and $sk$ the private key.
  • Enc: takes as input a public key $pk$ and a message $m$ from the plaintext space. It outputs a ciphertext $c\rightarrow\textsf{Enc}_{pk}(m)$.
  • Dec: takes as input a private key $sk$ and a ciphertext $c$, and outputs a message $m$ or a special symbol $\bot$ denoting failure.
It is required that for every $(pk,sk)$ and plaintext message $m$, it holds that $\textsf{Dec}_{sk}(\textsf{Enc}_{pk}(m))=m$. The IND-CPA security ensures the scheme leaks no useful information under a chosen plaintext attack.
IV Problem Settings
We now introduce Private Individual Computation, a new paradigm for privacy-preserving computation that offers several benefits: (1) It provides a formal privacy guarantee (in the differential privacy sense) for a wide range of data-driven computational tasks. (2) Since the computation is performed on user-sanitized data, it avoids the need for heavy cryptographic protocols, allowing for scalability to handle large datasets and user bases. (3) The privacy amplification effect results in significantly better utility compared to direct sanitization and computation using Local Differential Privacy (LDP). In this section, we will first present a few motivating applications and then formally define Private Individual Computation as an ideal functionality.
Figure 1: An illustration of taxi-hailing in the PIC model. Besides (sanitized) location information, each user also encapsulates a one-time random public key into the message to the shuffler.
IV-A Motivating Applications
We describe three prevalent exemplar computation tasks: spatial crowdsourcing, location-based social systems, and federated learning with incentives:
Spatial crowdsourcing. A spatial crowdsourcing system typically consists of three roles: users, workers, and the orchestrating server (the service platform). It proceeds through the following major steps:
  • I.
    Task submission: each user $i\in G_{a}$ submits a task (e.g., taxi calling requests, sensing/photo requests) $x_{i}=(i,l_{i},v_{i})$ containing the location $l_{i}$ and possibly other information $v_{i}$;
  • II.
    Worker reporting: every enrolled worker $j\in G_{b}$ reports $x_{j}$ containing the location $l_{j}$ and other information $v_{j}$;
  • III.
    Task assignment: the server receives $\{x_{i}\}_{i\in G_{a}}$ and $\{x_{j}\}_{k\in G_{b}}$, and outputs a matching $M:G_{a}\times G_{b}\mapsto\{0,1\}$ between $G_{a}$ and $G_{b}$ based on some criterion (e.g., minimizing total traveling costs, or maximizing matches).
  • IV.
    Task performing: users and workers retrieve matching results and collaboratively complete the task.
One example of taxi-haling is illustrated in Figure 1.
Location-based social systems. In a location-based social system, a set of users and a server interact as the following:
  • I.
    Querying: Each participating user $i\in G_{a}$ submits a request $x_{i}=(i,l_{i},v_{i})$, which includes information like location $l_{i}$ and preferences $v_{i}$.
  • II.
    Generating recommendation: The server receives $\{x_{i}\}_{i\in G_{a}}$ and generates a list of recommendations for each user based on specific criteria, such as proximity and user preferences.
  • III.
    Retrieving: Finally, each user retrieves the recommendation results.
Federated learning with incentives. Federated learning involves a set of users and a server:
  • I.
    Submitting gradient: Each user $i\in G_{a}$ in each epoch computes an intermediate gradient information $x_{i}\in[-1,1]^{d}$ with its local model and data, then submits to the server;
  • II.
    Computing incentive: The server computes the average gradient $\overline{x}=\frac{1}{n}\sum_{i\in[n]}x_{i}$. To incentive participation, the server may reward users with monetary tokens (e.g., via cryptocurrency) according to a profit allocation algorithm $V:[-1,1]^{d\times n}\times[-1,1]^{d}\mapsto[0,1]$ (e.g., Shapley value [71]).
  • II.
    Receiving incentive: Each user retrieves the token and claims its monetary incentive.
A fundamental distinction between the above applications and those currently studied in the shuffle model is that each participant now expects an output that differs individually, rather than a single collectively aggregated output. Another distinction is that a participant might need to securely communicate with other participants after receiving the individualized computation results, such as the matched user and worker in spatial crowdsourcing will communicate with each other to accomplish the task, and the matched users in location-based social systems would like to securely contact each other afterward. Informally, in such applications, it is necessary to safeguard data privacy so that, aside from the party who generates the data, no one can be certain about that party’s data (up to the leakage allowed by differential privacy). To amplify privacy, we also need to maintain anonymity so that for any given message, an adversary (such as the server, an observer, or an unmatched user) should only know that this message comes from a user within a particular group, but nothing more. Following the convention in the shuffle model, the shuffler is trusted to provide anonymity. The shuffler knows the random permutation used in the shuffling process and will not leak it to other parties, although the shuffler is prevented from observing plaintext messages through encryption.
IV-B The Ideal Functionality
Following the ideal-real world paradigm, we capture private individual computation formally as an ideal functionality $\mathcal{F}_{\scriptscriptstyle\textsf{PIC}}$, which is shown in Figure 2.
Functionality $\mathcal{F}_{\scriptscriptstyle\textsf{PIC}}$ Parameters: $m\in\mathbb{N}$; $m$ groups of parties $G_{1},G_{2},\cdots,G_{m}$, where $n_{i}=|G_{i}|$ denotes the number of parties in group $G_{i}$; the data randomization mechanisms $\mathcal{R}_{i}$ for group $G_{i}$; a server ${S}$. Functionality: Upon receiving $n=\sum_{i\in[m]}n_{i}$ inputs $\{x_{i,j}\}_{i\in[m],j\in[|G_{i}|]}$ from all users, and the description of a function $f$ to be computed over parties’ inputs from the server, do the following: Compute $x^{\prime}_{i,j}\leftarrow R_{i}(x_{i,j})$ for all $i\in[m]$ and $j\in[|G_{i}|]$. Sample $m$ random permutations $\pi_{1},\pi_{2},\cdots,\pi_{m}$, where $\pi_{i}:[n_{i}]\rightarrow[n_{i}]$, perform shuffle over inputs and obtain $L=\left(\{x_{1,\pi_{1}(j)}^{\prime}\}_{j\in[n_{1}]},\cdots,\{x_{m,\pi_{m}(j)}^{\prime}\}_{j\in[n_{m}]}\right)$, and compute ($\left(\{y_{1,\pi_{1}(j)}\}_{j\in[n_{1}]},\cdots,\{y_{m,\pi_{m}(j)}\}_{j\in[n_{m}]}\right)\leftarrow f(L)$. Send $y_{i,\pi_{i}(j)}$ to party $u_{i,j}$ for all $i\in[m]$ and $j\in n_{i}$. Additionally, send $(L,f(L))$ to $S$.
Figure 2: The functionality $\mathcal{F}_{\scriptscriptstyle\textsf{PIC}}$
Essentially, the ideal functionality represents a fully trusted party that interacts with $m$ groups of users and one server $S$. It sanitizes the input from each user, shuffles the inputs randomly, and then applies a function to compute the output for each user. In the end, each user receives their individual function output, while the server receives a list of the shuffled sanitized user data and a list of the function outputs. It captures the functional requirements of individual computation in the real world: a server performs a computing task using the joint inputs from a set of users (which are sanitized and shuffled), and each user receives an individualized output. It also captures the security requirements: each party receives precisely the specified output, and nothing more.
Remark 1 A careful reader may notice that the ideal functionality does not explicitly capture differential privacy. This omission is intentional for the sake of security analysis. In the security analysis, we decouple the privacy requirements into two sets of proofs: the first set demonstrates that our concrete protocol, when executed by real-world parties, realizes the ideal functionality. This means no additional information about parties’ inputs is revealed, except for the output given to each party. The second set of proofs establishes that the output given to each party conforms with differential privacy.
V A Concrete Protocol
V-A The Protocol
We now present a concrete protocol for PIC. Because there are already secure protocols for shuffling, to simplify the description, we present the protocol in the $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$-hybrid model, in which parties can communicate as usual, and in addition have access to an ideal functionality $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$ that does the shuffling. The ideal functionality $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$ (Figure 3) is parameterized with a list of parties that are corrupted by the adversary and collude with the server. For those parties, the adversary should know the correspondence between their messages before and after shuffling, hence $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$ leaks this part of the permutation to the adversary.
Functionality $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$ Parameters: $n\in\mathbb{N}$; $n$ parties $P_{1},P_{2},\cdots,P_{n}$; a server ${S}$; the corrupted party set $\mathcal{C}$; the leakage $\mathcal{L}(\pi)=\{(i,\pi(i))\}_{i\in[\mathcal{C}]}$ for the permutation $\pi$ being used. Functionality: Upon receiving $n$ inputs $\{x_{i}\}_{i\in[n]}$ from $P_{1},P_{2},\cdots,P_{n}$, respectively. Sample a random permutation $\pi\in\mathbf{S}_{n}$. Define $\{y_{i}\}_{i\in[n]}$ such that $y_{i}=x_{\pi(i)}$. Send $\{y_{i}\}_{i\in[n]}$ to the server $S$. Additionally if $S\in\mathcal{C}$, send $\mathcal{L}(\pi)$ to the adversary $S$.
Figure 3: The functionality $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$
The PIC protocol is outlined below:
  1. 1.
[…]
  2. 2.
    We denote the $j$-th user in group $G_{i}$ as $u_{i,j}$. Each user generates a key pair $(pk_{i,j},sk_{i,j})\leftarrow\textsf{Gen}(\lambda)$. Each user then randomizes their private data $x_{i,j}$ with mechanism $\mathcal{R}_{i}$ and obtains $x_{i,j}^{\prime}\leftarrow\mathcal{R}_{i}(x_{i,j})$. Then the sanitized input is concatenated with their own public key, and encrypted with the server’s public key $x_{i,j}^{\prime\prime}\leftarrow\textsf{Enc}_{pk_{c}}(pk_{i,j}||x^{\prime}_{i,j})$.
  3. 3.
    Each user in group $G_{i}$ invokes $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$ with $x_{i,j}^{\prime\prime}$, and $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$ outputs the shuffled messages $\{x_{i,\pi_{i}(j)}^{\prime\prime}\}_{j\in[n_{i}]}\leftarrow\mathcal{S}(\{x_{i,j}^{\prime\prime}\}_{j\in[n_{i}]})$ to the server, where $\pi^{-1}_{i}$ is the (secret) random permutation used during $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$ for group $G_{i}$.
  4. 4.
    The server decrypts each set of shuffled messages and obtains a list $L$ for all $m$ groups as:
    $L=$ ($\left(\{pk_{1,\pi_{1}(j)}||x_{1,\pi_{1}(j)}^{\prime}\}_{j\in[n_{1}]},\cdots\right.$
    $\ \ \ \left.\{pk_{m,\pi_{m}(j)}||x_{m,\pi_{m}(j)}^{\prime}\}_{j\in[n_{m}]}\right)$
    It then computes the function $f$ over $L$ to produce output for each anonymous user:
    ($\left(\{y_{1,\pi_{1}(j)}\}_{j\in[n_{1}]},\cdots,\{y_{m,\pi_{m}(j)}\}_{j\in[n_{m}]}\right)\leftarrow f(L).$
  5. 5.
    The server publishes the computation results to a public bulletin board as a list of pairs: $\{(pk_{i,\pi_{i}(j)},\textsf{Enc}_{pk_{i,\pi_{i}(j)}}(y_{i,\pi_{i}(j)})\}_{j\in[n_{i}]}$, for each group $i\in[m]$.
  6. 6.
    Every user downloads the list, finds in the list the entry with their own public key, and decrypts the payload to get the computation result.
Remark 2 In the final step, we adopt the simplest strategy for users to retrieve their results without the server knowing which entry belongs to whom. If bandwidth is a concern, it can be replaced by a more sophisticated (and computationally more expensive) Private Information Retrieval protocol (e.g. [46]).
Remark 3 To eventually accomplish the spatial crowdsourcing (e.g., taxi-hailing services) or location-based social system tasks, the computation result $y_{i,\pi_{i}(j)}$ will contain the matched users of the user $\pi_{i}(j)\in G_{i}$. That is, $y_{i,\pi_{i}(j)}$ will encapsulate a list of public keys and noisy location information about the matched users of the user $\pi_{i}(j)\in G_{i}$. After receiving the individualized computation results in the final step, each user $\pi_{i}(j)$ can then securely contact the matched users using the public keys in $y_{i,\pi_{i}(j)}$ (possibly with the help of a public communication channel, such as a public bulletin).
V-B Security Analysis
This section presents the formal security analysis of the protocol in the previous section. In the following, we will consider a semi-honest adversary who can statically corrupt parties in the protocol. That is, the adversary will faithfully follow the protocol specifications but try to learn more information than allowed through protocol interaction. Also, before running the protocol, the adversary specifies the corrupted parties. The adversary controls the corrupted parties and knows their internal states. We use $\mathcal{C}$ to denote the collection of corrupted parties and $\mathcal{C}\subset G_{1}\cup G_{2}\cdots G_{n}\cup\{S\}$.
We first show that our protocol securely realizes the ideal functionality $\mathcal{F}_{\scriptscriptstyle\textsf{PIC}}$ in the $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$-hybrid model. This means the protocol leaks no more information than what is allowed by $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$ and $\mathcal{F}_{\scriptscriptstyle\textsf{PIC}}$. More precisely, each user gets their output from $\mathcal{F}_{\scriptscriptstyle\textsf{PIC}}$, the server gets $(L,f(L))$, and in the case of colluding with some users in $G_{i}$, the partial permutation $\mathcal{L}(\pi_{i})$. Formally we have the following theorem:
Theorem V.1 (Security)
The Private Individual Computation (PIC) protocol in §V securely computes $\mathcal{F}_{\scriptscriptstyle\textsf{PIC}}$ in the $\mathcal{F}_{\scriptscriptstyle\textsf{Shuffle}}$-hybrid model in the presence of any PPT adversary with static corruption.
The proof can be found in Appendix -A. The proof is simulation-based. It shows that for any PPT adversary $\mathcal{A}$ in the real world, there exists a PPT simulator $\mathcal{S}$ in the ideal world that can generate a simulated view given the corrupted parties’ inputs and outputs. Security means that the simulated view is indistinguishable from the view of real-world execution.
The above theorem states that the adversary learns strictly no more than the allowed output and leakage by engaging in the protocol execution. Next, we will show how much differential privacy we can get in the presence of such an adversary with such knowledge. We consider an honest user $u_{\scriptscriptstyle i^{*},j^{*}}$, where $i^{*}\in[m],j^{*}\in[n_{i^{*}}]$. At the same time, the set of corrupted users in $G_{i^{*}}$ is denoted as $C_{i^{*}}\subset G_{i^{*}}$. Differential privacy in our case means that on two neighboring inputs $X=\big{(}X_{1},\cdots,X_{i^{*}}\cdots,X_{m}\big{)}$ and $X^{\prime}=\big{(}X_{1},\cdots,X^{\prime}_{i^{*}}\cdots,X_{m}\big{)}$, the output and leakage obtained by the adversary, denoted as $\mathcal{A}(X)$ and $\mathcal{A}(X^{\prime})$, are close in distribution. Formally, we have the following theorem:
Theorem V.2 (Differential Privacy)
The Private Individual Computation protocol satisfies $(\epsilon_{c},\delta)$-DP, i.e.
$\max(D_{\epsilon_{c}}(\mathcal{A}(X)||\mathcal{A}(X^{\prime})),D_{\epsilon_{c}}(\mathcal{A}(X^{\prime})||\mathcal{A}(X)))\leq\delta.$
In particular, when all users in $G_{i^{*}}$ use an identical $\epsilon$-LDP mechanism as the data randomizer $\mathcal{R}_{i^{*}}$, and for $n^{\prime}_{i^{*}}=|G_{i^{*}}-C_{i^{*}}|\geq 8(e^{\epsilon}+1)\log(2/\delta)$ we have:
$\epsilon_{c}=\log\Bigg{(}1+\frac{e^{\epsilon}-1}{e^{\epsilon}+1}\Big{(}\sqrt{\frac{32(e^{\epsilon}+1)\log 4/\delta}{n^{\prime}_{i^{*}}}}+\frac{4(e^{\epsilon}+1)}{n^{\prime}_{i^{*}}}\Big{)}\Bigg{)}.$ (1)
The analysis can be divided into two cases: in the first case, $S\in\mathcal{C}$, i.e. S is corrupted. In this case, since the honest user locally randomizes their input, the input enjoys at least $\epsilon$-DP. Then the shuffling will amplify the privacy guarantee. Since the corrupted server receives the partial permutation as the leakage, the amplification depends on the number of uncorrupted users in the same group ($n^{\prime}_{i^{*}}$) as the honest user. The amplified $\epsilon_{c}$ can then be derived following [34, 35]. In the second case where $S\not\in\mathcal{C}$, the knowledge of the adversary is $y_{i,j}$ for each corrupted user $u_{i,j}\in\mathcal{C}$. The tricky part is that how much information $y_{i,j}$ leaks depends on the function $f$ being evaluated by the server. Hence we consider the worst case where $f$ output $y_{i,j}=L$. Continuing along the same line of thought, we conclude that the level of differential privacy assurance is no less than in the first case. The full proof can be found in Appendix -B.
Remark 4 After decryption, the server obtains $pk_{i,j}||x^{\prime}_{i,j}$, where $pk_{i,j}$ is a public key not sanitized by the local randomizer. Despite the presence of the public key, $pk_{i,j}||x^{\prime}_{i,j}$ and $x^{\prime}_{i,j}$ are equivalent in terms of privacy amplification. This is because (1) the public key is random and generated independently of $x_{i,j}$, so prepending it to $x^{\prime}_{i,j}$ does not affect the local DP guarantee—it is the same as $x^{\prime}_{i,j}$ itself; and (2) all public keys follow an identical probability distribution across all users in the group $G_{i}$, ensuring that the privacy amplification effect via shuffling is not degraded.
V-C Discussion on Post-computation Communication
In certain scenarios, such as spatial crowdsourcing and location-based social systems, there may be additional user-to-user communications following the execution of the PIC protocol. For instance, consider a taxi-hailing application where passengers are in a group $G_{1}$ and taxi drivers in a group $G_{2}$. After receiving the matching result at the end of the protocol, the passengers must send their locations to the matched drivers, who need to know where to pick them up. On the other hand, the drivers also need to share their identities and locations with the matched passengers. Inevitably, a party has to sacrifice their privacy to the matched parties.
The post-computation communication may also have privacy implications for other users not in the matched pair. Privacy amplification via shuffling against an adversary relies on the number of parties that remain anonymous. Recall that in Equation 1, the amplified $\epsilon_{c}$ relies on the number $n^{\prime}_{i^{*}}=|G_{i^{*}}-C_{i^{*}}|$ of uncorrupted users in a particular group $G_{i^{*}}$. From the perspective of the adversary, if the post-computation communication compromises the anonymity of an additional set $U$ of the users in $G_{i^{*}}$, then $n^{\prime}_{i^{*}}$ becomes $|G_{i^{*}}-C_{i^{*}}-U|$. Accordingly, the privacy amplification effect for users in $G_{i^{*}}$ that are still anonymous is weakened.
We emphasize that preventing the loss or weakening of privacy via technical means is not feasible, because the information is necessary for the proper functioning of the application. However, managerial countermeasures, such as ensuring sufficiently large user groups and limiting post-computation exposure according to the need-to-know principle, can mitigate potential privacy risks arising from post-computation communication.
VI Optimal Randomizers
VI-A Inadequacy of Existing Randomizers
In PIC, each user first sanitizes their data using an $\epsilon$-LDP randomizer. The design of the randomizer significantly impacts the utility of the tasks. While the PIC model can be seen as an extension of the shuffle model, it has unique characteristics that render the existing LDP randomizers commonly used in the shuffle model inadequate.
The main discrepancy between the shuffle model and the PIC model is that the former emphasizes statistical utility, whereas the latter focuses on the utility of each report. The shuffle model aims to estimate certain statistics from the noisy data collected from users, so it cares about how close the estimation is to the true value of the desired statistic. In the literature, utility is often measured by the expected square error (i.e., the variance) bound of the estimation:
$\max_{\scriptscriptstyle T\in\mathbb{X}^{n}}\mathbb{E}[\|\tilde{f}(T)-f(T)\|_{2}^{2}]=\max_{\scriptscriptstyle T\in\mathbb{X}^{n}}\textsf{Var}[\tilde{f}(T)]$
where $f$ is a statistical function, and $\tilde{f}$ is its estimation output by the shuffle protocol. On the other hand, in the PIC model, the tasks are often non-statistical. For example, in location-based matching, the required computation is to take two users’ locations and compute the distance between them. Hence, the above utility measure is no longer suitable. It is more natural to measure utility by the single report error:
$\max_{x_{j}\in\mathbb{X}}\mathbb{E}[\|\mathcal{R}(x_{j})-x_{j}\|_{2}^{2}],$
where $\mathcal{R}$ is the LDP-randomizer employed by the users. Additionally, in the PIC model, user data is typically multi-dimensional (e.g., location, gradient), making sanitization significantly more difficult compared to scalar data. Another notable characteristic of the PIC model is that the local differential privacy budget $\epsilon$ is relatively large, often scaling linearly with $\log(n^{\prime}_{i^{*}})$, as implied by Theorem V.2. These factors together create issues when existing LDP randomizers are applied directly in the PIC model.
To understand the problem, we first examine a class of LDP randomizers [65, 86, 59, 51] that operate by sampling a few dimensions from $[d]$. Each user submits an incomplete report that contains only the sampled dimensions (with added noise) from their local data. On the positive side, this strategy reduces the amount of noise added to the sampled dimensions. On the negative side, the unsampled dimensions are missing. In statistical estimation tasks, the estimation is made using all reports, each covering some dimensions. Therefore, the incompleteness of a single report is less important, and better utility can be achieved. However, in the PIC model, where the focus shifts to the error of individual reports, this strategy may lead to worse results.
There have been LDP randomizers that submit complete reports. One obvious strategy is to explicitly split the local budget into $d$ parts and then apply a one-dimensional LDP mechanism independently to each dimension, or implicitly distribute the budget among dimensions, as seen in the Laplace [30], PlanarLaplace [2], PrivUnit [12], and PrivUnitG mechanisms [6]. However, these approaches are sub-optimal in the high budget regime. Specifically, even if we use the optimal one-dimension randomizer [7], splitting the budget across each dimension and then applying any randomizer for each dimension will result in a mean squared error (MSE) of at least $\frac{d}{(e^{\epsilon/d}-1)^{2/3}}$. The Laplace/PlanarLaplace mechanisms introduce an MSE rate of $\frac{d}{\epsilon^{2}}$, while the PrivUnit/PrivUnitG mechanisms incur an MSE rate of $\frac{d}{\min\{\epsilon,\epsilon^{2}\}}$. In contrast, later we will show that the MSE rate can be improved to $(e^{\epsilon}-1)^{-2/(d+2)}$ (see Section VI-C and Appendix -D).
Another strategy for submitting complete reports involves using additional randomization techniques such as random projection, data sketches, public randomness, or quantization. Randomizers employing this strategy [76, 36, 73, 17, 50, 4] avoid the issue of incomplete reports but introduce additional noise because of the extra randomization. While the additional noise is not significant in the low budget regime (i.e., $\epsilon=O(1)$) that is typical in the LDP model, it becomes dominant in the PIC model where the local budget can be as large as $\tilde{O}(\log n^{\prime}_{i^{*}})$. The resulting additional error will never diminish even when $\epsilon\to+\infty$.
VI-B Randomizer Design
We now introduce an asymptotically optimal randomizer, tailored for the PIC model. The randomizer uses an LDP mechanism, which we termed as Minkowski Response. For ease of presentation, here we will focus primarily on the $\ell_{2}$ case (and the $\ell_{+\infty}$ case in Appendix -D), but the mechanism can be generalized to other Minkowski distances.
Without loss of generality, we assume the user data domain to be an $\ell_{2}$-bounded hyperball $\mathbb{X}=\{x\ |\ x\in\mathbb{R}^{d}\ and\ \|x\|_{2}\leq 1\}$. Most, if not all, real-world data domains can be normalized to $\mathbb{X}$ (e.g. gradient vector, set-valued data, and location data). We also denote an $\ell_{2}$-bounded hyperball with radius $r$ centered at any $x\in\mathbb{R}^{d}$ as follows:
$\mathbb{B}_{r}(x)=\{x^{\prime}\ |\ x^{\prime}\in\mathbb{R}^{d}\ and\ \|x^{\prime}-x\|_{2}\leq r\},$
and it is shorted as $\mathbb{B}_{r}$ when $x=\vec{0}$.
Minkowski Response works by first defining a distance $r$ based on the local privacy budget as the following:
$r=\big{(}(e^{\epsilon}-1)^{1/(d+2)}-1\big{)}^{-1}.$
Then given the input domain $\mathbb{X}$, the output domain $\mathbb{Y}_{r}$ is defined by expanding $\mathbb{X}$ by $r$:
$\mathbb{Y}_{r}=\{y\ |\ y\in\mathbb{R}^{d}\ and\ \exists x\in\mathbb{X}\ that\ y\in\mathbb{B}_{r}(x)\}.$
Figure 4: The probability design of Minkowski response mechanism with a radius $r$. Illustrated are three inputs $x_{1},x_{2}$ and $x_{3}$, along with their respective cap areas.
For any input $x\in\mathbb{X}$, Minkowski Response outputs an output $y\in\mathbb{Y}_{r}$ with relatively high probability in the cap area $\mathbb{B}_{r}(x)$ and relatively low probability in remaining domain $\mathbb{Y}_{r}\backslash\mathbb{B}_{r}(x)$ (see Figure 4). Formally:
$y=\left\{\begin{array}[]{@{}ll@{}}\text{uniform}(\mathbb{B}_{r}(x)),&\text{with prob. $\frac{\mathrm{V}(\mathbb{B}_{r})\cdot(e^{\epsilon}-1)}{\mathrm{V}(\mathbb{Y}_{r})+\mathrm{V}(\mathbb{B}_{r})\cdot(e^{\epsilon}-1)}$;}\\ \text{uniform}(\mathbb{Y}_{r}),&\text{with prob. $\frac{\mathrm{V}(\mathbb{Y}_{r})}{\mathrm{V}(\mathbb{Y}_{r})+\mathrm{V}(\mathbb{B}_{r})\cdot(e^{\epsilon}-1)}$,}\\ \end{array}\right.$ (2)
where $V(*)$ denote the volume of the corresponding domain.
Lastly, $y$ is debiased to $\tilde{x}$ so that $\mathbb{E}[\tilde{x}]={x}$ as follows:
$\tilde{x}=y\cdot\frac{\mathrm{V}(\mathbb{Y}_{r})+\mathrm{V}(\mathbb{B}_{r})\cdot(e^{\epsilon}-1)}{V(\mathbb{B}_{r})\cdot(e^{\epsilon}-1)}.$ (3)
It is obvious that the output of Minkowski Response is informative in every dimension. Therefore it avoids problems brought up by incomplete reports. Also, intuitively it has a better utility because the mechanism is more likely to output a value near the true value (i.e., in $\mathbb{B}_{r}(x)$), than from other parts of the output domain (i.e. $\mathbb{Y}_{r}\backslash\mathbb{B}_{r}(x)$). When the budget $\epsilon$ gets large, the error rate of Minkowski Response decays faster than in previous LDP mechanisms. More specifically, the decay rate of Minkowski Response is $(e^{\epsilon}-1)^{-2/(d+2)}$ (Equation 6 in Appendix -D), while that of the previous mechanisms is $d/(e^{\epsilon/d}-1)^{2/3}$ or $d/\epsilon^{2}$. Therefore, the utility advantage of Minkowski Response becomes more significant when $\epsilon$ gets larger. When $n^{\prime}_{i^{*}}\to+\infty$ (and thus $\epsilon\to+\infty$), $r$ becomes 0, and the error goes to zero (i.e. no additional error).
VI-C Analysis of Minkowski Response
The local privacy guarantee of the randomizer is presented in Theorem VI.1.
Theorem VI.1 (Local Privacy Guarantee)
Given input domain $\mathbb{X}=\mathbb{B}_{1}$, the Minkowski response mechanism defined in Equation 2 satisfies $\epsilon$-LDP.
Proof:
It is observed that the output probability distribution in Equation 2 is valid for any input $x\in\mathbb{X}$, the probability density in the cap area $\mathbb{B}_{r}(x)$ is $\frac{e^{\epsilon}}{\mathrm{V}(\mathbb{Y}_{r})+\mathrm{V}(\mathbb{B}_{r})\cdot(e^{\epsilon}-1)}$, and the density in the non-cap area $\mathbb{Y}_{r}\backslash\mathbb{B}_{r}(x)$ is $\frac{1}{\mathrm{V}(\mathbb{Y}_{r})+\mathrm{V}(\mathbb{B}_{r})\cdot(e^{\epsilon}-1)}$. Therefore, for any $x,x^{\prime}\in\mathbb{X}$, we have $\frac{\mathbb{P}[\mathcal{R}(x)=y]}{\mathbb{P}[\mathcal{R}(x^{\prime})=y]}\leq e^{\epsilon}$ for all possible $y\in\mathbb{Y}_{r}$, establishing the local $\epsilon$-DP guarantee of the Minkowski response mechanism $\mathcal{R}$. ∎
Next, we analyze the utility of the Minkowski Response in the PIC model. As previously mentioned, in the PIC model, the single report error is a more appropriate measure of utility compared to the statistical errors used in the conventional shuffle model. In Theorem VI.2, we examine the single report error in the PIC model and establish its lower bound. The proof is provided in Appendix -C.
Theorem VI.2 (Error Lower Bounds)
Given $d\in\mathbb{N}$, $\epsilon_{c}>0$, $\delta\in(0,0.5]$, $\mathbb{X}=\mathbb{B}_{1}(\{0\}^{d})$, then for any randomizer $\mathcal{R}:\mathbb{X}\mapsto\mathbb{R}^{d^{\prime}}$ such that $\mathcal{S}\circ\mathcal{R}(X)$ and $\mathcal{S}\circ\mathcal{R}(X^{\prime})$ are $(\epsilon_{c},\delta)$-indistinguishable for all possible neighboring datasets $X,X^{\prime}\in\mathbb{X}^{n}$, and for any estimator $f:\mathbb{R}^{d^{\prime}}\mapsto\mathbb{R}^{d}$, we have:
$\max_{x\in\mathbb{X}}\mathbb{E}\big{[}\|f\circ\mathcal{R}(x)-x\|_{2}^{2}\big{]}\geq\tilde{\Omega}\big{(}{1}/{n^{\frac{2}{d+2}}}\big{)}.$
The above theorem suggests that in the PIC model, for any randomizer, the single report error is at least $\tilde{\Omega}\big{(}{1}/{n^{\frac{2}{d+2}}}\big{)}$. Clearly, a randomizer offers better utility if its error is closer to this bound. For the randomizer using Minkowski Response, as presented in section VI-B, we can demonstrate that its single report error has an upper bound. This is summarized in Theorem VI.3, with the proof provided in Appendix -D.
Theorem VI.3 (Error Upper Bounds)
For any $d\in\mathbb{N}$, $\epsilon_{c}>0$, $\delta\in(0,0.5]$, $\mathbb{X}=\mathbb{B}_{1}(\{0\}^{d})$, if $\epsilon_{c}\leq O(1)$ and $n>\max\{16\log(1/\delta),\frac{2^{d+7}\log(1/\delta)}{(e^{\epsilon_{c}}-1)^{2}}\}$, then there exist a randomizer $\mathcal{R}:\mathbb{X}\mapsto\mathbb{R}^{d}$ such that $\mathcal{S}\circ\mathcal{R}(X)$ and $\mathcal{S}\circ\mathcal{R}(X^{\prime})$ are $(\epsilon_{c},\delta)$-indistinguishable for all possible neighboring datasets $X,X^{\prime}\in\mathbb{X}^{n}$, and:
$\max_{x\in\mathbb{X}}\mathbb{E}\big{[}\|\mathcal{R}(x)-x\|_{2}^{2}\big{]}\leq O\Bigg{(}\Big{(}\frac{\log(1/\delta)}{n\epsilon_{c}^{2}}\Big{)}^{\frac{2}{d+2}}\Bigg{)}.$
We observe that in most applications, $(\epsilon_{c},\delta)$ are given as fixed system parameters. If we consider $(\epsilon_{c},\delta)$ as constants, then the error upper bound of the Minkowski Response randomizer in Theorem VI.3 is $\tilde{O}\left(\frac{1}{n^{2/(d+2)}}\right)$, which matches the error lower bound of the PIC model in Theorem VI.2. This implies that the utility of the Minkowski Response is asymptotically optimal. In contrast, using existing LDP randomizers in the PIC model results in a larger single report error of $\tilde{O}\left(\frac{d}{n^{2/(3d)}}\right)$ or $\tilde{O}\left(\frac{d}{\log^{2}n}\right)$ (note that $d>1$ and $n$ is often not small). Although asymptotic notations describe behavior as $n\rightarrow\infty$, in practice, the utility advantage of the Minkowski Response becomes noticeable without $n$ being very large: in our experiments, the Minkowski Response randomizer outperforms existing LDP randomizers in the PIC model once $n$ reaches the order of $10^{2}$. If we compare Minkowski Response in the PIC model to using LDP directly (without shuffling), the utility advantage is even greater: any randomizers in the LDP model must endure a single report error of $\Omega\left(\frac{d}{\epsilon_{c}^{2}}\right)$ when $\epsilon_{c}\leq O(1)$ [12, 28].
VII Experimental Evaluation
We evaluate the efficacy of our PIC protocol and Minkowski randomizer. We compare the utility of our proposal against state-of-the-art works in the context of three representative individual computation tasks: spatial crowdsourcing, location-based social systems, and federated learning with incentives.
VII-A Spatial Crowdsourcing
Datasets We use two real-world datasets: GMission dataset [18] for scientific simulation, and EverySender dataset [82] for campus-based micro-task completion. Details about the two datasets are summarized in Table II, including the number of users/workers, location domain range, and serving radius of workers about these two datasets (more information can be found in Appendix -F). We normalize location data to the domain $[-1,1]\times[-1,1]$ and scale the serving radius to $0.41.0\cdot\frac{1-(-1)}{5-0}=$ 0.4 correspondingly.
TABLE II: Summary statistics of spatial crowdsourcing datasets.
Dataset users workers location domain serving radius
GMission 713 532 $[0,5.0]\times[0,5.0]$ $1.0$
EverySender 4036 817 $[0,5.0]\times[0,5.0]$ $1.0$
LDP randomizers We use the Minkowski response for location randomization. As a comparison, in the local model of DP (e.g., in [85, 80, 84]), we compare with existing mechanisms including Laplace [31], Staircase [38], PlanarLaplace with geo-indistinguishability [2], SquareWave [58], PrivUnit [12] and its Gaussian variant PrivUnitG in [6].
Server-side algorithms Two commonly used server-side algorithms are evaluated as concrete tasks: minimum weighted full matching [52] and maximum matching [47]. The two algorithms have different optimization objectives, thus later we will show the results for task-specific utility for each of them in addition to single report errors. The minimum weighted full matching aims to minimize the overall traveling costs between users and workers. The maximum matching aims to maximize the number of successfully matched user/worker pairs, where users outside the workers’ serving radius ($0.4$) are deemed unreachable.
Experimental results We first present the single report errors, quantified by the expected $\ell_{2}$ distance between the reported location and the true location. In Fig. 5 (a) and (b), we compare the single report errors within the LDP model. The Minkowski randomizer’s error is on par with other state-of-the-art mechanisms when $\epsilon\leq 1.0$ and significantly lower when $\epsilon\geq 2.0$. Fig. 5 (c) and (d) illustrate the single report errors in the PIC model. For both datasets, the Minkowski randomizer performs better, with its utility advantage increasing as the global privacy budget $\epsilon_{c}$ grows. Additionally, it is evident that the error is generally higher for all randomizers in the LDP model compared to the PIC model. This discrepancy is due to the lack of privacy amplification in the LDP model, highlighting the benefits of employing the PIC model.
Figure 5: Expected $\ell_{2}$ distances of reported locations to true locations on GMission and EverySender dataset.
Next, we compare the utility when using the minimum weight matching algorithm, as shown in Fig. 6. Task-specific utility is evaluated by the total travel costs, which is the sum of the actual Euclidean distances between all matched users/workers:
$\sum\nolimits_{(i,j)\in G_{a}\times G_{b}}\llbracket M(i,j)>0\rrbracket\cdot\|l_{i}-l_{j}\|_{2},$
[…]
Figure 6: Travel costs of minimum weighted matching on GMission and EverySender dataset.
Finally, we present the utility comparison using the maximum matching algorithm, as shown in Fig. 7. In this case, task-specific utility is assessed by the successful matching ratio:
$\frac{\sum_{(i,j)\in G_{a}\times G_{b}}\llbracket M(i,j)>0\rrbracket\cdot\llbracket\|l_{i}-l_{j}\|_{2}\leq\tau\rrbracket}{\min\{|G_{a}|,|G_{b}|\}}.$
For both datasets, the matching ratio over clear data is 100%. The figure demonstrates that the PIC model enhances the matching ratio for all randomizers due to privacy amplification effects. The Minkowski randomizer in the PIC model significantly outperforms the others and the matching ratio approaches an acceptable level for practical use with a reasonable degree of privacy protection.
Figure 7: Success ratios of maximum matching in spatial crowdsourcing on GMission and EverySender dataset.
VII-B Location-based Social Systems
TABLE III: Details of location social network datasets.
Dataset check-ins location domain
Gowalla (San Francisco) 138368 $[37.54,37.79]$$\times$$[-122.51,-122.38]$
Foursquare (New York) 227428 $[40.55,40.99]$$\times$$[-74.27,-73.68]$
Datasets We use two real-world datasets: Gowalla dataset [21] and Foursquare dataset [89]. Gowalla and Foursquare are location-based social network websites where users share their locations by checking-in. Details about the two datasets are summarized in Table III. As before, the location data is normalized to $[-1,1]\times[-1,1]$.
LDP randomizers The LDP randomizers used are the same as those in Section VII-A.
Server-side algorithm The server performs the radius-based nearest neighbor (NN) search for the users, which is a common task in location-based social networks [21]. We set the search radius $\tau$ to 0.2, so that each user has several hundreds or thousands of neighbors, varying due to check-in densities. Note that in this application scenario, the returned neighbors may be deanonymized in the post-computation phase. Hence, the actual privacy guarantee in the PIC model depends on the number of users who remain anonymous (see discussion in Section V-C). Considering this, and each user normally has no more than $n\cdot\frac{\pi\tau^{2}}{2^{2}}<n\cdot 3.2\%$ neighbors, we use privacy amplification population $\lfloor n\cdot 90\%\rfloor$ instead of the group size $n$ when calculating the local budget given the global $\epsilon_{c}$.
Figure 8: Expected $\ell_{2}$ distances of reported locations in location-based social systems with the local model of DP.
Figure 9: Expected $\ell_{2}$ distances of reported locations in location-based social systems with PIC model.
Experimental results We first present the single report errors in the LDP model (Fig. 8) and the PIC model (Fig. 9). In our experiments, each check-in report is treated as if it were submitted by a separate user, with all users in the same group. We tested two scenarios: one with a random subset of 10,000 check-ins and the other using all check-ins. On both datasets, PIC demonstrates better utility than LDP, and the Minkowski randomizer consistently performs the best across all settings. Additionally, it is evident that the number of anonymous users is a crucial parameter, significantly impacting utility.
Figure 10: F1 scores of nearest neighbor queries (LDP model).
Figure 11: F1 scores of nearest neighbor queries (PIC model).
The task-specific utility metric we use for nearest neighbor queries is the $F_{1}$ score. Let $N_{i}$ denote the set of true neighbors of user $i$ within an $\ell_{2}$-distance $\tau$, and let $\widehat{N}_{i}$ denote the retrieved neighbor set computed using the sanitized reports. The precision of nearest neighbor queries is defined as $\frac{\sum_{i\in[n]}|N_{i}\cap\widehat{N}_{i}|}{\sum_{i\in[n]}|\widehat{N}_{i}|}$, and the recall is defined as $\frac{\sum_{i\in[n]}|N_{i}\cap\widehat{N}_{i}|}{\sum_{i\in[n]}|N_{i}|}$. The $F_{1}$ score is then calculated as:
$F_{1}\text{ score}=2\cdot\frac{\text{precision}\cdot\text{recall}}{\text{precision}+\text{recall}}.$
In the LDP model (Fig. 10), the $F_{1}$ score is low, even with a large $\epsilon$. While the group size affects the number of neighbors each user has, it does not impact the $F_{1}$ score. Conversely, in the PIC model (Fig. 11), the $F_{1}$ score is significantly higher. With a large group, the Minkowski randomizer can achieve an $F_{1}$ score of 0.9 with a stringent global budget $\epsilon_{c}=1$ and nearly 1 if the budget is loosened to $\epsilon_{c}=3$.
VII-C Federated Learning with Incentives
In this application, the users collaboratively train a model by Federated learning, and the server decides each user’s incentive by how much their local gradient contributes to the global model. To deliver monetary incentive rewards, we assume the use of an untraceable cryptocurrency (e.g., ZCash [47]). Each user can generate an additional public/privacy key pair according to the specification of the cryptocurrency, derive a wallet address from the public key, and append the wallet address to the end of their report. Based on the Shapley values, the server determines the monetary incentives and distributes them to the users using appended wallet addresses.
Datasets We utilize the MNIST handwritten digit dataset to train a simple neural network using federated learning. The MNIST dataset comprises 60,000 images, with 50,000 designated as training samples. Each user receives one training sample and trains a neural network model, as detailed in Table IV, which contains $d=4292$ trainable parameters. In each round, $s=10,000$ users are randomly selected. These selected users locally compute the gradient vector using Stochastic Gradient Descent (SGD), then subsample 0.15% of the gradient vector dimensions (with unsampled dimensions set to zero), and clip the gradient values at a threshold of $c=0.00015$. Each user submits a sanitized version of the subsampled, clipped gradient vector as their report.
TABLE IV: Model architecture of the neural network.
Layer Parameters
Convolution 8 filters of 4$\times$4, stride 2
Max-pooling 2$\times$2
Convolution 6 filters of 5$\times$5, stride 2
Max-pooling 2$\times$2
Softmax 10 units
LDP randomizers We use all LDP randomizers as before except for PlanarLaplace, which is unsuitable for multi-dimensional gradient vectors. The randomizers sanitize the non-zero values in the subsampled gradient vectors while leaving the zero values unaffected.
Server-side algorithm The server performs the following steps: first, it takes a sanitized gradient vector and uses randomizer-specific algorithms to estimate the true values of the sampled dimensions. These vectors with estimated values are denoted as $\hat{g}_{i}$. The server then aggregates the $\hat{g}_{i}$ vectors into a global gradient vector using a simple sum-and-average method. This global gradient vector is published so that users can update their local models. Additionally, the server computes the Shapley value for each user, which measures the marginal contribution of each $\hat{g}_{i}$ (see Appendix -H for details).
Experimental results We train each model for 80 rounds using Federated Learning. The utility comparison results are summarized in Tables V and VI, each reflecting a different global privacy budget. In these tables, we present the single report utility (Gradient $\ell_{2}$ error) and the task-specific utility (Shapley $\ell_{2}$ error). The Gradient $\ell_{2}$ error is calculated using the estimated gradients $\hat{g}_{i}$ and the true gradients $g_{i}$ (clipped and with all unsampled dimensions set to 0): $\frac{\sum_{i\in S}\|\hat{g}_{i}-{g}_{i}\|_{2}}{|S|}$, where $S$ is the set of randomly selected users. The Shapley $\ell_{2}$ error is computed as follows:$\frac{\sum\nolimits_{i\in S}\|\widehat{\texttt{Shapley}}_{i}-{\texttt{Shapley}}_{i}\|_{2}}{|S|}$, where $\widehat{\texttt{Shapley}}_{i}$ is derived from $\hat{g}_{i}$ and ${\texttt{Shapley}}_{i}$ is obtained from the unsanitized $g_{i}$. For reference, we also include the final models’ accuracy in the tables.
TABLE V: Utility comparison of federated learning with incentives: global privacy budget $(1,0.01/50000)$.
Setting
Randomization
Mechanism
Test
Accuracy
Gradient
$\ell_{2}$ Error
Shapley
$\ell_{2}$ Error
local model Staircase [38] $18.29\%$ $15.30$ $0.0586$
Squarewave [58] $13.91\%$ $13.23$ $0.0590$
PrivHS [29] $28.17\%$ $2.44$ $0.0587$
Laplace [31] $21.22\%$ $2.53$ $0.0575$
PrinUnit [12] $23.28\%$ $2.37$ $0.0592$
Minkowski $18.10\%$ $14.45$ $0.0571$
PIC model Staircase [38] $32.51\%$ $0.781$ $0.0563$
Squarewave [58] $33.75\%$ $0.604$ $0.0585$
PrivHS [29] $65.47\%$ $0.267$ $0.0428$
Laplace [31] $72.6\%$ $0.154$ $0.0508$
PrinUnit [12] $74.02\%$ $0.157$ $0.0438$
Minkowski $\mathbf{78.68\%}$ $\mathbf{0.093}$ $\mathbf{0.0363}$
TABLE VI: Utility comparison of federated learning with incentives: global privacy budget $(3,0.01/50000)$.
Setting
Randomization
Mechanism
Test
Accuracy
Gradient
$\ell_{2}$ Error
Shapley
$\ell_{2}$ Error
local model Staircase [38] $19.68\%$ $5.09$ $0.0519$
Squarewave [58] $18.24\%$ $4.34$ $0.0574$
PrivHS [29] $41.52\%$ $0.850$ $0.0561$
Laplace [31] $28.64\%$ $0.845$ $0.0575$
PrinUnit [12] $39.02\%$ $0.803$ $0.0537$
Minkowski $20.73\%$ $3.731$ $0.0545$
PIC model Staircase [38] $44.58\%$ $0.545$ $0.0563$
Squarewave [58] $53.02\%$ $0.407$ $0.0552$
PrivHS [29] $74.49\%$ $0.190$ $0.0383$
Laplace [31] $74.86\%$ $0.104$ $0.0335$
PrinUnit [12] $77.42\%$ $0.098$ $0.0253$
Minkowski $\mathbf{83.43\%}$ $\mathbf{0.055}$ $\mathbf{0.0219}$
We observe in the tables that the utility in the LDP model is significantly worse than in the PIC model, as expected. This is consistent across all metrics and the final model accuracy. In the local model, the performance of Staircase, Squarewave, and Minkowski mechanisms is poorer compared to the others, but for different reasons. Staircase and Squarewave are designed for single-dimensional data, requiring the privacy budget to be split across dimensions when sanitizing vectors, which leads to reduced utility. For Minkowski, the issue lies in the 80-round federated learning process, where the privacy budget for each round is relatively small (e.g., between 0.2 and 0.75). As previously mentioned, the Minkowski mechanism is intended to achieve better utility with a large local privacy budget. In small privacy budget scenarios, it offers no advantage and may even perform worse. In the PIC model, the performance of Staircase and Squarewave remains poor. However, Minkowski now performs the best among all randomizers. This improvement is due to the privacy amplification effect in PIC, which increases the local privacy budget for each round to between 4.0 and 6.0.
VIII Conclusion
Privacy-preserving computation with Differential Privacy (DP) holds great potential for leveraging personal information. While the shuffle model offers a rigorous DP guarantee with enhanced utility, its application is confined to statistical tasks. In this paper, we introduce a novel paradigm called Private Individual Computation (PIC), which extends the shuffle model to scenarios where each user requires personalized outputs from the computation. We demonstrate that PIC can be realized using an efficient protocol that relies on minimal cryptographic operations while maintaining the advantages of privacy amplification through shuffling. To further enhance utility, we developed a local randomizer specifically designed for PIC. We provide formal proofs of the protocol’s security and privacy, as well as the asymptotic optimality of the randomizer. Extensive experiments validate the superiority of the PIC protocol and the randomizer, showcasing their performance across three major application scenarios and various real-world datasets.